### v1

# Disentangled Representation with Cross Experts Covariance Loss for Multi-Domain Recommendation

Abstract.
Multi-domain learning (MDL) has emerged as a prominent research area aimed at enhancing the quality of personalized services. The key challenge in MDL lies in striking a balance between learning commonalities across domains while preserving the distinct characteristics of each domain. However, this gives rise to a challenging dilemma. On one hand, a model needs to leverage domain-specific modules, such as experts or embeddings, to preserve the uniqueness of each domain. On the other hand, due to the long-tailed distributions observed in real-world domains, some tail domains may lack sufficient samples to fully learn their corresponding modules. Unfortunately, existing approaches have not adequately addressed this dilemma. To address this issue, we propose a novel model called Crocodile, which stands for Cross-experts Covariance Loss for Disentangled Learning. Crocodile adopts a multi-embedding paradigm to facilitate model learning and employs a Covariance Loss on these embeddings to disentangle them. This disentanglement enables the model to capture diverse user interests across domains effectively. Additionally, we introduce a novel gating mechanism to further enhance the capabilities of Crocodile. Through empirical analysis, we demonstrate that our proposed method successfully resolves these two challenges and outperforms all state-of-the-art methods on publicly available datasets. We firmly believe that the analytical perspectives and design concept of disentanglement presented in our work can pave the way for future research in the field of MDL.
Multi-domain Learning; Disentangled Representation Learning; Dimensional collapse
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Crocodile: Cross Experts Covariance for Disentangled Learning in Multi-Domain Recommendation; October 21-25, 2024; Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
((a)) Dillema of Interest Diversity v.s. Collapse Prevention
((b)) gAUC in Kuairand1k and AliCCP
Figure 1. The proposed Crocodile successfully resolves the dilemma of interest diversity v.s. collapse prevention and achieves the best performance regarding gAUC.
Online recommendation systems (Broder, 2008) are now essential tools for businesses to effectively reach their target audiences, servicing billions of internet users daily. Empowered by advanced algorithms and vast traffic, these systems play a crucial role in modeling user experiences and generating revenue for advertisers. However, traditional methods still face substantial challenges due to the dynamic nature of online platforms, which are marked by diverse user preferences across domains. To address these challenges, researchers have increasingly turned to multi-domain learning (MDL) (Chang et al., 2023; Zhou et al., 2023; Sheng et al., 2021; Yang et al., 2022; Li et al., 2023) as a promising approach to enhance the performance by involving information from multiple related domains.
Unfortunately, MDL for recommendation often confronts the challenge of preference conflict, where a user might exhibit divergent behaviors across domains. To remedy this, the MDL community aims to develop domain-specific components, including STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and PEPNet (Chang et al., 2023) have progressively integrated domain-specific components.
Interestingly, we observe a trend in MDL methods towards the development of bottom-level domain-specific modules. In this work, we uncover the mechanisms behind this trend. To the best of our knowledge, this is the first study to reveal the mechanism of increasing interest diversity to resolve preference conflict. Furthermore, we propose a novel metric, the Diversity Index (DI), to quantify interest diversity, as shown in Fig 1(a).
Based on our discovery, we notice that domain-specific embedding (Su et al., 2023; Guo et al., 2023a) incurs the highest diversity and hence should be able to resolve preference conflict. However, naively deploying domain-specific embeddings does not turn out to be fruitful, we observe negative transfer in domains with fewer data. Delving further, we find out the core issue is sub-optimal embeddings for the domain with less data, i.e., they are dimensionally collapsed. [^1] We adopt the note of dimensional collapse in the work of Chi et al. (2022) and utilize information abundance (IA) (Guo et al., 2023b) to quantify dimensional collapse.
As shown in Fig. 1(a), our analysis demonstrates that existing works struggle to maintain these two properties: maintaining diversity and avoiding dimensional collapse, thus resulting in the sub-optimal performance shown in Fig. 1(b). As a result, it remains an open question whether we can incorporate both properties.
To this end, we propose a novel a novel Cross-experts Covariance Loss for Disentangled Learning model (Crocodile) framework, which we build this Covariance Loss (CovLoss) to alleviate embedding collapse while in the meantime enabling learning diverse user interests across various domains. To further unleash the power of Crocodile, we develop a novel Prior Informed element-wise Gating (PeG) mechanism to work with CovLoss.
Empirical analysis on two public datasets confirms that Crocodile achieves state-of-the-art results. Further analysis also demonstrates that Crocodile can simultaneously mitigate embedding collapse and learn diverse representations. Ablation experiments reveal that each proposed component contributes positively to the overall performance. Our contributions can be briefly summarized as follows:
  • We discover the mechanism of previous works in maintaining diversity, especially the trend of focusing more on bottom-level modules. We propose a novel metric, Diversity Index (DI), to quantify the diversity and conclude that the trend of previous works is optimizing to increase DI.
  • We reveal that existing works fall into the dilemma of maintaining diversity and avoiding collapse, hence resulting in inferior performance.
  • We propose the Crocodile approach, which successfully balances both properties. Empirical evidence on two public datasets demonstrates that Crocodile achieves state-of-the-art performance and resolves the aforementioned dilemma through comprehensive analysis.
2. Related Work
2.1. Multi-domain Recommendation
Multi-domain learning (MDL) is a promising topic that has garnered attention from industrial recommendations. Many MDL methods have made great efforts to model complex multi-domain interests. Among them, MoE-based methods are widely applied in industrial recommendation systems, such as Taobao (Sheng et al., 2021) , Meituan (Chang et al., 2023) and Kuaishou (Zhou et al., 2023).
STAR (Sheng et al., 2021) utilizes the dot product of domain-specific and domain-shared weight matrices to capture domain-specific interests. HiNet (Zhou et al., 2023) proposes two expert-like components to capture cross-domain correlations and within-domain information explicitly. PEPNet (Chang et al., 2023) introduces PPNet and EPNet, which incorporate domain information through gating. These methods often innovate from the perspective of making more bottom-level modules be domain-personalized within the model.
Recently, STEM (Su et al., 2023) proposed a multi-embedding (ME) structure in multi-task learning (MTL) and step forward in allocating more bottom-level domain-personalized structure, i.e., embedding. It achieved better results compared to both single and multi-embedding versions of MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) in MTL experiments. However, STEM fails in the multi-domain recommendation. We will discuss the reasons in detail in the subsequent sections.
Besides, some other paradigms are applied for multi-domain recommendation. AdaSparse (Yang et al., 2022) and ADL (Li et al., 2023) applied adaptive learning mechanism. EDDA (Ning et al., 2023a), COAST (Zhao et al., 2023) and UniCDR(Cao et al., 2023) provided graph learning-based solutions.
2.2. Disentangled Representation Learning
In the realm of disentangled representation learning (DRL), methods based on spatial or statistical relationships have been proposed.
On the one hand, methods based on spatial relationships achieve disentangling by manipulating the distances or angles between representations. OLE (Lezama et al., 2018), dot product(Wu et al., 2021) and cosine similarity(Liu et al., 2021) push representations towards orthogonal. GAN-based models (Chen et al., 2016; Yu and Welch, 2021; Khrulkov et al., 2021) promote the discrimination among representations. However, they are computationally expensive and not conducive to online recommendation.
On the other hand, statistical methods view representations as instances drawn from high-dimensional distributions and design loss based on statistical measures. Distance correlation (dCorr)(Liu et al., 2022a) encourages lower dCorr between different factors. ETL (Chen et al., 2023) models the joint distribution of user behaviors across domains with equivalent transformation. VAE-based methods (Ma et al., 2019; Wang et al., 2022; Liu et al., 2022b) encourage higher KL divergence between dimensions. However, they often assume a normal prior distribution, which is challenging to satisfy in the context of sparse interactions (Rendle, 2010), especially under sample volume imbalance of multi-domain recommendation.
Besides, some methods integrate both aspects mentioned above. VICReg (Bardes et al., 2021; Shwartz-Ziv et al., 2024) employs covariance and variance as loss components within representations to promote independence across dimensions and prevent collapse while utilizing Euclidean distance between representations to ensure semantic similarity among representations of different patches of an image.
2.3. Dimensional Collapse
Dimensional collapse(Chi et al., 2022) is defined as representations that end up spanning a lower-dimensional space, which is evident through the pronounced imbalance in the distribution of singular values. Chi et al. (2022) found that SMoE structure (Fedus et al., 2022; Lepikhin et al., 2020) may lead to dimensional collapse due to the influence of gating. The work of (Guo et al., 2023b) claimed that this problem is also observed in popular models (e.g., DCN V2 (Wang et al., 2021) and Final MLP (Mao et al., 2023)). To measure the collapse, they proposed the Information Abundance (IA) (Guo et al., 2023b) metric, defined as the ratio between the sum and the maximum l1-norm value of the target singular values. A small IA indicates severe insignificance of tail dimensions and potential dimensional collapse.
Beyond a single perspective, some methods (Hua et al., 2021) built the bridge between representation disentanglement and dimensional collapse prevention. It also claimed feature decorrelation helps alleviate dimensional collapse by reducing strong correlations between axes and promoting more diverse and informative representations.
3. On the Challenges of Multi-Domain Recommendation
As we have discussed in Sec. 1, existing works fall into the dilemma of avoiding dimensional collapse and maintaining interest diversity. In this section, we will discuss these two properties thoroughly.
3.1. Preventing Dimensional Collapse
According to Fig 1(a), domain-specific embeddings demonstrate its effectiveness in capturing diverse user interests. However, the naive adoption of domain-specific embeddings does not yield the expected outcomes. In fact, the methods with domain-specific embeddings, i.e., ME-PLE (Tang et al., 2020; Su et al., 2023) and STEM (Su et al., 2023), both performance deteriorate in the domain with less data. For example, in domain S6 of Kuairand1k (Gao et al., 2022) dataset, the ME-PLE and STEM show -0.4% and -0.1% lower AUC than the single domain method. This phenomenon has sparked our deep investigation into its underlying mechanism.
Intuitively, the domain-specific embeddings are trained solely on data from the corresponding domain, which results in insufficient samples for domains with less data. Insufficient training data fails to provide adequate variance for high-dimensional representations, resulting in high-dimensional representations residing within a lower-dimensional subspace, called dimensional collapse (Guo et al., 2023b; Chi et al., 2022; Hua et al., 2021).
To measure the collapse, we adopt the IA metric (Guo et al., 2023b), which is defined as $IA(E)=||\sigma||_{1}/||\sigma||_{\infty}$, where $E$ is the embedding of feature (e.g., item id) and $||\cdot||_{1},||\cdot||_{\infty}$ represent the $l1$-norm and $\infty$-norm, respectively.
Remark 1.
Intuitively, singular values $\sigma_{i}$ represent the significance of $i$-th singular value, and IA captures the balance of the singular value distribution. Thus, a small IA indicates severe insignificance of tail dimensions and potential dimension collapse.
((a)) $log(IA)$ of Experts
((b)) $log(IA)$ of Item Embeddings
Figure 2. Information Abundance ($log(IA)$) dynamics of STEM’s bottom layer of experts and item embeddings on Kuairand1k. High and Low denote the items with the highest and lowest frequencies, respectively.
Taking STEM as the example of methods applying domain-specific embedding, we analyzed its training IA dynamics of experts and embeddings in Fig. 2. In Kuairand1k, S1 and S6 are the domains with the most and least data. In Fig. 2(a), the S6-specific expert exhibits a lower IA compared to S1, indicating that domains with a smaller data volume are more susceptible to dimensional collapse at the expert level. In Fig. 2(b), we further categorized items into five groups based on their exposure frequency and observed the collapse of item ID embeddings for both the highest and lowest frequency groups. We discovered that domain S6 exhibits a lower IA than the overall S6, indicating a more severe collapse, regardless of the group. Within each domain, the low-exposure group consistently shows a lower IA than the high-exposure group. This further corroborates the notion that a lower exposure count (i.e., data volume) is more prone to dimensional collapse.
Both analytical and empirical evidence suggest that inadequate data of domain can lead to dimensional collapse of the domain-specific embedding, hence leading to inferior performance.
3.2. Interest Diversity
Intuitively, the ability to capture various user interests across domains can be directly characterized by the diversity of representations. When facing conflict preference across domains, a set of diverse representations is enabled to encode both prefer/non-prefer semantics and thus provide distinct vectors to towers of those conflict domains.
To this end, we propose a novel metric, the Diversity Index (DI), to quantify the representations’ ability to capture conflict preference across domains. Specifically, DI depicts the ratio of samples that contain both large and small norm of representations within the whole conflict samples, which is formulated as [^2] Here $AF$ and $SF$ are abbreviations for all forward and specific forward respectively, aligned with the concepts in  (Su et al., 2023).:
$\mathcal{D}_{con(i,j)}$ contains preference conflict samples between domain $i$ and $j$. $\mathbb{I}[\text{{condition}}]$ is the indication function, which is assigned to 1 if the condition is true; otherwise, 0. $M$ represents experts, $\mathbf{O}_{k}^{(p)}$ means the output of $p$-th expert and $k$-th sample, $\tau_{t}$ and $\tau_{b}$ are the thresholds of big and small l2-norm.
Remark 2.
As shown in Fig. 1(a), ME-MMoE (Guo et al., 2023a) exhibits significantly lower diversity compared to STEM (Su et al., 2023), and its improvement in diversity is relatively limited when compared to MMoE (Ma et al., 2018b). This suggests that the representations learned by ME-MMoE (Guo et al., 2023a) lack sufficient ability to solve preference conflict challenges.
Figure 3. Crocodile Architecture, which consists of a Multi-Embedding (ME) layer, a Cross-expert Covariance Loss (CovLoss), and a Prior informed Element-wise Gating (PEG) mechanism.
4. Methodology
We propose a multi-embedding architecture with Cross-experts Covariance Loss for Disentangled Learning (Crocodile) to resolve the dilemma, containing Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG), as shown in Fig. 3.
4.1. Multi-Embedding Layer
The raw input features $x=\{x_{1},x_{2},\cdots,x_{F}\}$ contains $F$ fields. We employ a multi-embedding (ME) architecture (Su et al., 2023), which utilizes $M$ sets of embedding lookup tables. Each embedding looked up from each respective table is fed into the corresponding expert. Thus, in ME architecture, there are $M$ sets of experts. For a given $p$-th embedding table corresponding to $k$-th sample and $f$-th field, the looked-up embedding is represented as $e_{k,f}^{(p)}$. Thus, each $p$-th embeddings corresponded to $k$-th sample is:
(3) $e_{k}^{(p)}=[e_{k,0}^{(p)},e_{k,1}^{(p)},\cdots,e_{k,F}^{(p)}].$
4.2. Experts and Towers
Methods based on the Mixture of Experts (MoE) have been widely applied for multi-task (Ma et al., 2018b; Tang et al., 2020; Su et al., 2023) or multi-domain (Zhou et al., 2023; Chang et al., 2023) recommendation. Thus, we followed the same architecture and employed several sets of Multi-Layer Perceptrons (MLPs) as the experts and towers. Besides, we utilized a gating mechanism to compute the weights for the representation output from each expert. The output of the $p$-th expert can be represented as :
(4) $\mathbf{O}_{k}^{(p)}=MLP^{(p)}(e_{k}^{(p)}),$
where $\mathbf{O}_{k}^{(p)}\in\mathcal{R}^{1\times d}$ and ReLU serves as the inter-layer activation function. Through weighted by gating, the outputs $\mathbf{O}_{k}^{(\cdot)}$ of each expert are aggregated into a vector, denoted as $t_{k}^{(s)}\in\mathcal{R}^{1\times d}$, the input of $s$-th domain. The detail of gating will be introduced in Section 4.4.
Finally, each domain-specific tower outputs the final predicted value for the corresponding domain as :
(5) $\hat{y}_{k}^{(s)}=\sigma(MLP(t_{k}^{(s)})),$
where $\hat{y}_{k}^{(s)}$ is the predicted value of $k$-th sample in the $s$-th domain, $\sigma(\cdot)$ is the sigmoid function.
4.3. Cross-expert Covariance Loss
Many pioneering work (Shwartz-Ziv et al., 2024; Bardes et al., 2021; Ermolov et al., 2021) have highlighted the efficiency of tailoring the covariance into objective functions. However, they aimed to decorrelate the dimensions within each representation rather than disentangle among representations of all experts. To learn diverse user interests at the expert level, we propose the cross-expert covariance loss (CovLoss) to explicitly disentangle representations among experts, defined as:
(6) $\mathcal{L}_{Cov}=\frac{1}{d^{2}}\sum\limits_{p,q\in M\times M,p>q}||[\mathbf{O}^{(p)}-\overline{\mathbf{O}}^{(p)}]^{T}[\mathbf{O}^{(q)}-\overline{\mathbf{O}}^{(q)}]||_{1},$
where $\mathbf{O}^{(p)},\mathbf{O}^{(q)}\in\mathcal{R}^{N\times d}$ are the outputs of $p$-th and $q$-th experts, $\overline{\mathbf{O}}^{(p)}\in\mathcal{R}^{1\times d}$ is the average value of each dimension across all samples, and $||\cdot||_{1}$ is the $l1$-norm.
Finally, the overall loss function is defined as :
(7) $\mathcal{L}=\sum_{s=1}^{S}[\frac{1}{N_{s}}\sum_{k=1}^{N_{s}}\mathcal{L}_{BCE}(\hat{y}_{k}^{(s)},y_{k}^{(s)})]+\alpha\mathcal{L}_{Cov}$
where $N_{s}$ is the number of $s$-th domain’s samples, $\hat{y}_{k}^{(s)}$ is the label of $k$-th sample in domain $s$, and $\alpha$ is the weight of CovLoss.
4.4. Prior Informed Element-wise Gating Mechanism
CovLoss aims to disentangle different representations by minimizing the covariance between dimensions among different experts. However, the existing gating structure hinders the effectiveness of CovLoss. MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020)’s gating mechanism assigns an identical weight to all dimensions within each expert; the dimensions within each expert are mutually constrained. PEPNet (Chang et al., 2023)’s proposed Gate Neaural Unit operates at the dimension level as well. However, it is designed to scale each dimension separately within the representation but not to disentangle these representations. Consequently, we propose an Element-wise Gating mechanism (EG), which implements weight control over corresponding dimensions between experts, as shown in Fig. 3:
(8) $t_{k}^{s}$ $=g^{s}_{k}\odot\mathbf{O}_{k},$
where $\mathbf{O}_{k}\in\mathcal{R}^{K\times d}$ is the concatnation across all experts respected to $k$-th sample, and $\odot$ is the element-wise multiplication.
Moreover, we are also concerned about the parameter burden of ME structure. In the original ME-MMoE structure, the number of embeddings is tied to the number of domains. This is because the gating is performed using $g^{s}_{k}=Softmax(e_{k}^{s}W_{g}^{s})$, which means the embedding $e_{k}^{s}$ is still related to a $s$-th gating. Thus, when it is deployed in situations with a large amount of domains, the scale of parameters becomes unmanageable. To address this issue, we proposed Prior Informed Gating (PG), which involves an independent set of domain-unrelated prior embeddings $r_{k}$ to control the gating. $r_{k}$ takes the prior features as input, denoted as $x^{prior}$. For a given $k$-th sample and $f$-th field, the looked-up prior embedding is $e_{k,f}^{prior}$. Thus, the input of each gating is:
(9) $r_{k}=[e_{k,0}^{prior},e_{k,1}^{prior},\cdots,e_{k,F^{\prime}}^{prior}],$
where prior embedding contains $F^{\prime}$ fields. we will select user ID, item ID, and domain ID as the prior embedding fields. We defined the PG mechanism as:
(10) $g^{s}_{k}$ $=Softmax(r_{k}W_{g}^{s}),$
where $r_{k}\in\mathcal{R}^{1\times l}$ is the prior embedding of $k$-th sample, $W_{g}^{s}\in\mathcal{R}^{l\times K\times d}$ and $Softmax(\cdot)$ is operated across the second dimension.
Furthermore, PG can be understood as an attention (Vaswani et al., 2017)-like mechanism, comprising query, key, and value components corresponding to Prior embedding $r_{k}$, gating weight matrix $W_{g}^{s}$ of $s$-th domain, and expert outputs $\mathbf{O}_{k}$, respectively. Its significance lies in weighting the learnt user interests through the interaction of user, item, and domain information provided by query and key.
4.5. Understanding the CovLoss
CovLoss decorates the representations of different experts in an element-wise fashion. Beyond statistic understanding, CovLoss also directly addresses the core objective of cross-domain modeling: enhance the diversity of representations.
For simplification, we consider the model involving only two experts, outputting $\mathbf{x}$ and $\mathbf{y}$ drawn from two high-dimensional distributions $\mathcal{X}$ and $\mathcal{Y}$, respectively, and the $i$-th element of $\mathbf{x}$ and $\mathbf{y}$ are $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, respectively. Following (Shwartz-Ziv et al., 2024), we assume the randomness comes from the input of the network.
To maximize the diversity of the outputs from both experts, the objective for each pair of samples can be expressed as the expectation of l2-norm difference:
(11) $\mathrm{E}[(||\mathbf{x}||_{2}$ $-||\mathbf{y}||_{2})^{2}]=\mathrm{E}[||\mathbf{x}||_{2}^{2}+||\mathbf{y}||_{2}^{2}-2||\mathbf{x}||_{2}||\mathbf{y}||_{2}]$
$=\sum_{i=1}^{d}[\mu_{xi}^{2}+\sigma_{xi}^{2}]+\sum_{j=1}^{d}[\mu_{yj}^{2}+\sigma_{yj}^{2}]-2\mathrm{E}[\sqrt{\sum_{j\geq i}(x_{i}y_{j})^{2}}]$
$\geq-2\mathrm{E}[\sqrt{\sum_{i=1}^{d}\sum_{j=1}^{d}(x_{i}y_{j})^{2}}]$
$\geq-2\mathrm{E}[\sum_{i=1}^{d}\sum_{j=1}^{d}\sqrt{(x_{i}y_{j})^{2}}]=-2\sum_{j\geq i}\mathrm{E}[x_{i}y_{j}]$
$=-2\sum_{i=1}^{d}\sum_{j=1}^{d}[\mu_{xi}\mu_{y_{j}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}cov(x_{i},y_{j})}],$
where $x_{i},y_{j}\geq 0$ because the outputs from experts have been activated by ReLU before. According to the above Eq., minimizing $cov(x_{i},y_{j})$ can maximize the lower bound of diversity objective [^3] Although $\mu_{xi}\mu_{y_{j}}+cov(x_{i},y_{j})$ simultaneously incorporates both the first and second moments, they are nearly independent of each other in optimization.
5. Experiment
In this section, we provide comprehensive analyses to address the primary research questions (RQs) outlined below:
  • RQ1: How does the Crocodile perform compared with state-of-the-art recommenders in public datasets?
  • RQ2: Can Crocodile capture diverse user interests without introducing dimensional collapse issue?
  • RQ3: How does each proposed component contribute to performance gain?
  • RQ4: How sensitive is Crocodile to critical hyper-parameters and model capacity?
We employed two metrics to measure the model performance: AUC and gAUC. AUC measures the overall ranking performance, and gAUC evaluates the quality of intra-user item ranking.
5.1. Datasets
In this study, to examine the performance of the model in realistic recommendation, we carefully selected the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets for our experiments. For Kuairand1k, we chose five domains with top-5 data volumes. For AliCCP, we selected all three domains. Tab. 1 is the statistics for the two datasets after low-frequency filtering. We removed features with fewer than 10 exposures and replaced them with default values.
Table 1. Statistics of Kuairand1k and AliCCP datasets.
#Sample #User #Item #Impr in Domain
Kuairand1k (Gao et al., 2022) 11,648,743 1,000 188,148 2.4M(S0) / 7.8M(S1) / 0.4M(S2) / 0.9M(S4) / 0.2M(S6)
AliCCP (Ma et al., 2018a) 85,316,519 241,080 499,321 32M(S1) / 52M(S2) / 0.6M(S3)
5.2. Baselines
To establish a performance benchmark for comparison, we will select comparative methods from both MTL and MDL for multi-domain recommendation. We apply the SharedBottom (Caruana, 1997) method in single/multi-domain contexts and subsequently introduce popular MTL methods such as MMoE (Ma et al., 2018b), PLE (Tang et al., 2020) (both Single-Embedding and Multi-Embedding (Su et al., 2023) versions), and STEM (Su et al., 2023). Besides, we choose representative MDL methods, including STAR (Sheng et al., 2021), PEPNet (Chang et al., 2023), HiNet (Zhou et al., 2023), and AdaSparse (Yang et al., 2022).
5.3. Performance Evaluation (RQ1 and RQ2)
Table 2. Performance on Kuairand1k dataset. Bold and underline highlight the best and second-best results, respectively. * indicates that the performance difference against the second-best result is statistically significant at p-value¡0.05.
Methods AUC gAUC
S0 S1 S2 S4 S6 Overall S0 S1 S2 S4 S6 Overall
Single Domain 0.73272 0.74202 0.80179 0.72505 0.75559 - 0.60336 0.59610 0.54581 0.55085 0.60562 -
MTL (SE) for MDR Shared Bottom 0.73535 0.74271 0.82033 0.72552 0.75945 0.78574 0.60335 0.59694 0.56337 0.56299 0.62500 0.66087
MMoE 0.73639 0.74279 0.82090 0.72793 0.76005 0.78594 0.60188 0.59658 0.56278 0.56325 0.62721 0.66057
PLE 0.73716 0.74271 0.81759 0.72737 0.75807 0.78565 0.60585 0.59710 0.56234 0.56276 0.62908 0.66093
MTL (ME) for MDR ME-MMoE 0.73842 0.74281 0.81897 0.72637 0.75985 0.78586 0.60649 0.59874 0.56620 0.56080 0.63001 0.66188
ME-PLE 0.73645 0.74240 0.80785 0.72727 0.75138 0.78499 0.61084 0.59717 0.55982 0.56170 0.61744 0.66075
STEM 0.73741 0.74267 0.80526 0.72776 0.75449 0.78527 0.60820 0.59701 0.56250 0.56365 0.62595 0.66096
MDL PEPNET 0.73732 0.74275 0.81396 0.72767 0.76239 0.78574 0.60812 0.59589 0.55435 0.55931 0.62733 0.66018
HiNet 0.73285 0.74264 0.81955 0.72698 0.76451 0.78563 0.60303 0.59830 0.56416 0.56267 0.62790 0.66131
STAR 0.71890 0.74071 0.81586 0.72203 0.75954 0.78336 0.58028 0.59396 0.55229 0.55896 0.62419 0.65652
AdaSparse 0.73445 0.74235 0.81575 0.71845 0.76151 0.78481 0.60653 0.59729 0.56316 0.56128 0.63150 0.66031
Crocodile 0.73874 0.74430* 0.81646 0.72838 0.75817 0.78683* 0.61075 0.60068* 0.56964 0.56366 0.63579 0.66373*
[…]
Table 3. Performance on AliCCP dataset. bold and underline highlight the best and second-best results, respectively. * indicates that the performance difference against the second-best result is statistically significant at p-value¡0.05.
Methods AUC gAUC
S1 S2 S3 Overall S1 S2 S3 Overall
Single Domain 0.61299 0.56544 0.61735 - 0.58077 0.53023 0.58739 -
MTL (SE) for MDR Shared Bottom 0.61925 0.59355 0.61952 0.61885 0.58908 0.57252 0.59076 0.58982
MMoE 0.61951 0.59324 0.61969 0.61935 0.58907 0.57269 0.59065 0.58992
PLE 0.62152 0.59476 0.61949 0.61931 0.58911 0.57256 0.59070 0.58947
MTL (ME) for MDR ME-MMoE 0.62271 0.59741 0.62179 0.62131 0.59002 0.57258 0.59148 0.59048
ME-PLE 0.62352 0.58727 0.62063 0.62132 0.58974 0.55777 0.59130 0.59049
STEM 0.61978 0.58187 0.61978 0.61907 0.58941 0.55101 0.59132 0.59036
MDL PEPNET 0.62273 0.59239 0.61988 0.62034 0.58851 0.56080 0.59085 0.58975
HiNet 0.62042 0.59515 0.62003 0.62022 0.58881 0.57191 0.59015 0.58970
STAR 0.62039 0.59204 0.61886 0.61949 0.58705 0.56527 0.58932 0.58855
AdaSparse 0.62192 0.59656 0.62173 0.62189 0.58819 0.57170 0.58972 0.58924
Crocodile 0.62448 0.59951 0.62192 0.62233* 0.59039 0.57085 0.59204* 0.59123*
We compared Crocodile with state-of-the-art recommenders in the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets. All experiments are repeated 6 times and averaged results are reported, as shown in Tab. 2 and Tab. 3, respectively.
5.3.1. Overall Performance.
Our proposed Crocodile abrains the best overall AUC and gAUC in two datasets. In the Kuairand1k dataset, Crocodile outperforms the second-best method significantly by 0.09% and 0.19% in terms of overall AUC and gAUC, respectively. In the AliCCP dataset, Crocodile surpasses the second-best method significantly by AUC of 0.04% and gAUC of 0.07% than the second-best method.
Crocodile achieves larger AUC and gAUC almost for almost all domains in two datasets significantly compared to single domain method, with p-value¡0.05. Such positive transfer can be attributed to Crocodile’s better usage of multi-domain information and multi-embedding to enhance performance in each domain, without raising dimensional collapse.
Figure 4. Comparison of the singular value spectrum $log(\sigma)$ of item ID and user ID embeddings in Kuairand1k dataset. We reported $log(\sigma)$ of ME-PLE and STEM S6-specific embedding, while the original $log(\sigma)$ of single embedding or the average of other ME methods.
5.3.2. Comparison with MTL for MDR methods.
Multi-task learning (MTL) methods for Multi-Domain Recommendation (MDR) can be primarily categorized into single-embedding (SE) and multi-embedding (ME) paradigms (Su et al., 2023). We observed that some methods allocating domain-specific embedding underperform compared to SE methods, despite the fact they can capture diverse user interests in Fig. 5(a). For instance, in the comparison between PLE and ME-PLE on the Kuairand1k dataset in, Tab. 2, PLE achieved an overall AUC of $0.78565$ and gAUC of $0.66093$. In contrast, ME-PLE’s AUC and gAUC were $0.78499$ and $0.66075$, respectively, both lower than those of PLE.
To investigate the underlying reason, we conducted a domain-by-domain performance comparison and found that ME-PLE exhibited notably poorer results in the smallest domain, S6, with a negative transfer in AUC of -0.4% compared to the single domain. A similar negative transfer was also observed in STEM. We analyzed their embeddings via singular spectrum analysis (Guo et al., 2023b) and show the results in Fig. 4. We discovered that ME-PLE and STEM, in comparison to other methods, exhibited a clear dominance of top factors with much higher singular value over tail factors in the S6 embeddings, indicating the dimensional collapse issue.
In contrast, Crocodile’s embeddings exhibit a more balanced importance of singular values, thereby better supporting the high-dimensional representational space. Consequently, within domain S6, Crocodile outperforms ME-PLE by 0.68% in AUC and 1.8% in gAUC, and surpasses STEM by 0.37% in AUC and 0.98% in gAUC. Even though STEM learns the most diverse user interest, it still fails to surpass Crocodile due to dimensional collapse.
((a)) Crocodile vs MDL and MTL
((b)) Crocodile vs other Losses
Figure 5. Diversity Index (DI) of representations in the Kuairand1k dataset. There are five domains in Kuairand1k, resulting in 19 unique combinations. We plot the DIs on all combinations in descending order. We compared Crocodile with MDL and MTL methods in (a) and other disentanglement losses in (b).
5.3.3. Comparison with MDL methods.
The Multi-domain Learning (MDL) approaches typically evolve on a trend of progressively allocating domain-personalized components from top to bottom to capture more diverse user interest.
To empirically support this, we analyze the Diversity Index (DI) of both MDL and MTL methods in Fig. 5(a). Firstly, along with the trend from top to bottom, the DI of STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and STEM (Su et al., 2023) reveals an increased trend. Secondly, methods with an ME structure typically exhibit higher DI than their SE versions, e.g., ME-PLE has higher DI than PLE.
However, as discussed in Section 5.3.2, the allocation of domain-specific embeddings can lead to the side effect of dimensional collapse, which in turn may result in a negative transfer. Consequently, we are curious as to whether methods that do not employ domain-specific embeddings can achieve favorable outcomes.
For example, HiNet (Zhou et al., 2023) only allocates domain-specific components at the expert level and displayed the best ability to prevent dimensional collapse in Fig. 4. It thus got the best AUC of 0.76451 in domain S6, in Kuairand1k. However, HiNet still falls in suboptimal performance in overall AUC and gAUC in both datasets compared to Crocodile. This is because of its low ability to capture diverse user interests in  5(a).
Conversely, Crocodile achieves the second-best DI, only slightly trailing behind STEM, which employs domain-specific embeddings but faces the issue of dimensional collapse. Consequently, Crocodile attains the best overall AUC and gAUC.
5.3.4. Comparison with Other Forms of Losses.
Through comparison with MTL and MDL approaches, we found that Crocodile captures the diverse interests of users (Fig. 5(b)) without introducing the issue of dimensional collapse (Fig. 4), thereby achieving the best overall AUC and gAUC scores in Kuairand1k and AliCCP.
Thus, can other disentangled losses also achieve favorable diversity and corresponding positive outcomes compared to those without them? We compare CovLoss with several other disentangling losses, including Dot Product (Dot) (Wu et al., 2021), Cosine Similarity (Cos) (Liu et al., 2021), Distance Correlation (dCorr) (Liu et al., 2022a) and OLE (Lezama et al., 2018) (F-norm form).
We first analyzed the ability to capture user diverse interests in Fig. 5(b). We observed that all disentangled losses could achieve higher DI than ME-MMoE-PeG, where ME-MMoE-PEG is the basic structure without any disentangled loss compared to others.
Table 4. Comparison of different disentanglement losses with the base model on overall AUC and gAUC in Kuairand1k. * indicates that the performance difference against base is significant at p-value¡0.05.
Losses AUC Lift(%) gAUC Lift(%)
Base 0.78611 - 0.66208 -
+Dot (Wu et al., 2021) 0.78667 0.06% 0.66285 0.08%
+Cos (Liu et al., 2021) 0.78634 0.02% 0.66263 0.06%
+dCorr (Liu et al., 2022a) 0.78621 0.01% 0.66214 0.01%
+OLE (Lezama et al., 2018) 0.78650 0.04% 0.66319 0.11%
+Importance (Shazeer et al., 2017) 0.78632 0.02% 0.66208 0.00%
+Trans5 (Chen et al., 2023) 0.78550 -0.06% 0.66017 -0.19%
+CovLoss
0.78683 0.07% 0.66373 0.17%
Among them, we found that dCorr (Liu et al., 2022a) loss achieves the smallest improvement of only 0.01% and 0.01% for AUC and gAUC lift, respectively in Tab. 4. This is coherent with its related poor diversity compared to other disentangled losses in Fig. 5(b). In contrast, ultilizing CovLoss achieves the best AUC and gAUC with the lift of 0.07% and 0.17% compared to the Base model. This evidence supports that the ability to learn diverse interests is important to improve performance.
Furthermore, we are interested in exploring whether the transfer of other types of loss functions, e.g., load balance (Shazeer et al., 2017) and transfer learning (Chen et al., 2023), could yield performance enhancements. As shown in Tab. 4, the Importance loss (Shazeer et al., 2017) resulted in a limited improvement, whereas Trans5, one of the effective versions of equivalent transformation losses in (Chen et al., 2023), had a detrimental effect. This can be attributed to that these loss functions do not explicitly promote diversity in learning. Instead, they focus on balancing the contributions of various experts or enhancing the robustness of representations, which does not meet the expected outcomes.
5.4. Ablation Study (RQ3)
Table 5. Ablation study in Kuairand1k and AliCCP. ME and SE are shorts for multi- and single-embedding, respectively. B and B+C are shorts for BCE Loss and BCE Loss with CovLoss, respectively. * indicates that the performance difference against Base is significant at p-value¡0.05.
Methods Components Kuairand1k Performance AliCCP Performance
Embedding Gating Loss AUC Lift (%) gAUC Lift (%) AUC Lift (%) gAUC Lift (%)
Base (MMoE) SE - B 0.78594 - 0.66057 - 0.61935 - 0.58992 -
Ablation Method ME - B 0.78586 -0.01% 0.66188 0.13% 0.62131 0.20% 0.59048 0.06%
ME PG B 0.78585 -0.01% 0.66208 0.15% 0.62170 0.24% 0.59080 0.09%
ME PEG B 0.78611 0.02% 0.66208 0.15% 0.62227 0.29% 0.59117 0.12%
ME PG B+C 0.78642 0.05% 0.66333 0.28% 0.62214 0.28% 0.59124 0.13%
Crocodile ME PEG B+C 0.78683* 0.09% 0.66373* 0.32% 0.62237* 0.30% 0.59134* 0.14%
In this section, we empirically test the effectiveness of our proposed components with an ablation study. We eliminated CovLoss, PeG, and ME one by one and compared their performance on Kuairand1k and AliCCP, as shown in Tab 5.
All three components are critical since removing either of them leads to a significant performance drop on both datasets. In particular, CovLoss is more critical in the Kuairand1k dataset, while CovLoss and EG are equally important in AliCCP.
Table 6. gAUC metric along with different numbers of embeddings (#Emb) on ME-MMoE-PG and Crocodile. * indicates that the performance difference against ME-MMoE-PG is significant at p-value¡0.05.
#Emb 2 3 4 5 6 7 8 9 10
ME-MMoE-PG 0.66138 0.66157 0.66206 0.66232 0.66203 0.66196 0.66286 0.66279 0.66285
Crocodile 0.66198* 0.66284* 0.66324* 0.66373* 0.6637* 0.66394* 0.66395* 0.66432* 0.66397*
Lift (%) 0.06% 0.13% 0.12% 0.14% 0.17% 0.20% 0.11% 0.15% 0.11%
5.5. Hyper-Parameters and Embedding Sets (RQ4)
In this section, we investigate the sensitivity of the model’s performance to hyperparameters and the number of embeddings. For hyperparameters, the primary parameter is the weight of CovLoss in Eq. 7, i.e., $\alpha$. Taking Kuairand1k as an example, we find that as long as $\alpha$ is not too small (below $2\times 10^{-5}$), Crocodile’s AUC and gAUC remain ¿0.78590 and ¿0.66174, thus maintaining its status as the second-best method.
Additionally, we analyzed the impact of model capacity on performance by adjusting the number of embedding sets. At the minimum number of MEs (i.e., 2-sets), Crocodile achieves a gAUC of 0.66198 and an AUC of 0.78595, which still surpasses other MTL and MDL methods, even when the MTL(ME) methods apply for 5-sets of embeddings. Moreover, as the number of embeddings increases, the performance gains of Crocodile over ME-MMoE-PG are robust ($+0.06\%\sim+0.20\%$), as illustrated in Tab. 6.
5.6. Computational Complexity
The Multi-Embedding (ME) paradigm has already been widely adopted in industry (Zhaok et al., 2021; Ning et al., 2023b; Su et al., 2023; Guo et al., 2023b; Pan et al., 2024). Compared to these ME architectures, our proposed Crocodile introduces the CovLoss upon it, which only influences the training process while not the inference at all. Its complexity is not only small and negligible when compared with the ME-based backbone, but it can be further reduced by sampling.
The complexity of CovLoss is $\frac{M(M-1)}{2}\times(d^{2}N)$, shown lower complexity than dCorr (Liu et al., 2022a), which is $\frac{M(M-1)}{2}\times(d^{3}N+d^{2}N)$, where $M$ and $N$ are the number of experts and samples, and $d$ is the dimension of expert output. Besides, the complexity of experts in MMoE structure is $M\times\sum_{i=1}^{l-1}Nd_{i}d_{i-1}$, where $d_{i}$ is the hidden layer dimension for i-th layer in all $l$ layers. This is much higher than CovLoss in practice.
Moreover, we can reduce the complexity of calculating covariance by sampling. We discovered that selecting as low as 32 samples yields no significant difference in computed covariance norms (p-value¿ 0.1) in the testing dataset, contributing as much as 99.2% complexity reducing, as we originally selected $N=1024$.
6. Conclusion
Multi-domain learning is vital in enhancing personalized recommendations. Existing MTL methods fail to learn both dimensional-robust representations and diverse user interests. We propose Crocodile, which employs a Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG) based on the multi-embedding (ME) paradigm. Comprehensive experimental evaluations validate the effectiveness of Crocodile.

### v2

@@
-Disentangled Representation with Cross Experts Covariance Loss for Multi-Domain Recommendation
+Crocodile: Cross Experts Covariance for Disentangled Learning in Multi-Domain Recommendation
 Abstract.
-Multi-domain learning (MDL) has emerged as a prominent research area aimed at enhancing the quality of personalized services.
-The key challenge in MDL lies in striking a balance between learning commonalities across domains while preserving the distinct characteristics of each domain.
-However, this gives rise to a challenging dilemma.
-On one hand, a model needs to leverage domain-specific modules, such as experts or embeddings, to preserve the uniqueness of each domain.
-On the other hand, due to the long-tailed distributions observed in real-world domains, some tail domains may lack sufficient samples to fully learn their corresponding modules.
-Unfortunately, existing approaches have not adequately addressed this dilemma.
-To address this issue, we propose a novel model called Crocodile, which stands for Cross-experts Covariance Loss for Disentangled Learning.
-Crocodile adopts a multi-embedding paradigm to facilitate model learning and employs a Covariance Loss on these embeddings to disentangle them.
-This disentanglement enables the model to capture diverse user interests across domains effectively.
-Additionally, we introduce a novel gating mechanism to further enhance the capabilities of Crocodile.
-Through empirical analysis, we demonstrate that our proposed method successfully resolves these two challenges and outperforms all state-of-the-art methods on publicly available datasets.
-We firmly believe that the analytical perspectives and design concept of disentanglement presented in our work can pave the way for future research in the field of MDL.
+Multi-domain learning (MDL) has become a prominent topic in enhancing the quality of personalized services.
+It’s critical to learn commonalities between domains and preserve the distinct characteristics of each domain.
+However, this leads to a challenging dilemma in MDL.
+On the one hand, a model needs to leverage domain-aware modules such as experts or embeddings to preserve each domain’s distinctiveness.
+On the other hand, real-world datasets often exhibit long-tailed distributions across domains, where some domains may lack sufficient samples to effectively train their specific modules.
+Unfortunately, nearly all existing work falls short of resolving this dilemma.
+To this end, we propose a novel Cross-experts Covariance Loss for Disentangled Learning model (Crocodile), which employs multiple embedding tables to make the model domain-aware at the embeddings which consist most parameters in the model, and a covariance loss upon these embeddings to disentangle them, enabling the model to capture diverse user interests among domains.
+Empirical analysis demonstrates that our method successfully addresses both challenges and outperforms all state-of-the-art methods on public datasets.
+During online A/B testing in Tencent’s advertising platform, Crocodile achieves 0.72% CTR lift and 0.73% GMV lift on a primary advertising scenario.
 Multi-domain Learning; Disentangled Representation Learning; Dimensional collapse
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Crocodile: Cross Experts Covariance for Disentangled Learning in Multi-Domain Recommendation; October 21-25, 2024; Boise, Idaho, USAisbn: 978-1-4503-XXXX-X/18/06
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
-((a)) Dillema of Interest Diversity v.s. Collapse Prevention
+((a)) Dillema of Preserving Distinctiveness v.s. Sufficient Parameters Learning
 ((b)) gAUC in Kuairand1k and AliCCP
-Figure 1. The proposed Crocodile successfully resolves the dilemma of interest diversity v.s. collapse prevention and achieves the best performance regarding gAUC.
+Figure 1.
+Our proposed Crocodile successfully resolves the dilemma of preserving domain distinctiveness (measured by Diversity Index) v.s.
+sufficient parameters learning (measured by Information Abundance), and achieves the best performance measured by gAUC.
 Online recommendation systems (Broder, 2008) are now essential tools for businesses to effectively reach their target audiences, servicing billions of internet users daily.
@@
 However, traditional methods still face substantial challenges due to the dynamic nature of online platforms, which are marked by diverse user preferences across domains.
-To address these challenges, researchers have increasingly turned to multi-domain learning (MDL) (Chang et al., 2023; Zhou et al., 2023; Sheng et al., 2021; Yang et al., 2022; Li et al., 2023) as a promising approach to enhance the performance by involving information from multiple related domains.
-Unfortunately, MDL for recommendation often confronts the challenge of preference conflict, where a user might exhibit divergent behaviors across domains.
-To remedy this, the MDL community aims to develop domain-specific components, including STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and PEPNet (Chang et al., 2023) have progressively integrated domain-specific components.
-Interestingly, we observe a trend in MDL methods towards the development of bottom-level domain-specific modules.
-In this work, we uncover the mechanisms behind this trend.
-To the best of our knowledge, this is the first study to reveal the mechanism of increasing interest diversity to resolve preference conflict.
-Furthermore, we propose a novel metric, the Diversity Index (DI), to quantify interest diversity, as shown in Fig 1(a).
-Based on our discovery, we notice that domain-specific embedding (Su et al., 2023; Guo et al., 2023a) incurs the highest diversity and hence should be able to resolve preference conflict.
-However, naively deploying domain-specific embeddings does not turn out to be fruitful, we observe negative transfer in domains with fewer data.
-Delving further, we find out the core issue is sub-optimal embeddings for the domain with less data, i.e., they are dimensionally collapsed.
-[^1] We adopt the note of dimensional collapse in the work of Chi et al.
-(2022) and utilize information abundance (IA) (Guo et al., 2023b) to quantify dimensional collapse.
-As shown in Fig.
-1(a), our analysis demonstrates that existing works struggle to maintain these two properties: maintaining diversity and avoiding dimensional collapse, thus resulting in the sub-optimal performance shown in Fig.
+To address these challenges, researchers have increasingly turned to multi-domain learning (MDL) (Chang et al., 2023; Zhou et al., 2023; Sheng et al., 2021; Yang et al., 2022; Li et al., 2023b) as a promising approach to enhance the performance by involving information from multiple related domains.
+However, MDL faces conflicts of interest among users across different domains, which limits the potential for further exploitation of multi-domain information.
+Consequently, existing MDL models attempt to maintain domain distinctiveness through the implementation of domain-aware components along the upside-down trend, i.e., from Tower to Embedding.
+Specifically, Shared-Bottom (Caruana, 1997) and STAR (Sheng et al., 2021) focuses on making the tower domain-specific, while HiNet (Zhou et al., 2023), M-scan (Zhu et al., 2024) and HierRec (Gao et al., 2024) shift the focus to the expert level.
+PEPNet (Chang et al., 2023) follows this trend further by emphasizing domain-specific transformed embeddings.
+However, due to the domination of embeddings regarding the number of parameters in recommendation (Pan et al., 2024), these experts or transformation networks have limited capacities, making their ability to preserve each domain’s distinctiveness in doubt (refer to Finding 1 in Sec.
+2.1).
+Meanwhile, recent work in recommendation shows the efficacy of learning multiple separate embeddings for each feature.
+In particular, the Multi-Embedding paradigm (Guo et al., 2023a; Pan et al., 2024) was proposed in single-task learning for feature interaction models, aiming to resolve the dimensional collapse of embeddings caused by the Interaction-Collapse theory (Guo et al., 2023a).
+Regarding multi-task learning, a variant of multiple embeddings, namely the Shared and Task-specific EMbedding (STEM) (Su et al., 2023) was proposed to retain each task’s distinct characteristics via task-specific embeddings.
+Following these works, we try incorporating Shared and Domain-specific EMbeddings (SDEM) in multi-domain learning.
+However, these models are still not performing well.
+After careful analysis, we found that the embeddings are insufficiently learned, especially for those small domains (refer to Finding 2 in Sec.
+2.2).
+These small domains have much fewer training samples, making it challenging to learn independent embedding tables for them.
+In this paper, we propose a novel Diverse Multi-Embeddings paradigm, which adopts multiple shared embeddings with a novel disentanglement loss.
+Every embedding table is shared and hence optimized by all domains, overcoming the above-mentioned challenge of embedding insufficient learning, while the disentanglement loss promotes the diversity of these embeddings to preserve the underlying characteristics of all domains, as shown in Fig.
+1(a).
+Upon these multiple disentangled embeddings, a novel Prior Informed element-wise Gating (PeG) mechanism was proposed to route between these embeddings and domain-specific classification towers.
+We name our method Crocodile, standing for Cross-experts Covariance loss for Disentangled Learning.
+Empirical analysis conducted on two public datasets demonstrates that Crocodile achieves state-of-the-art results, as briefly shown in Fig.
 1(b).
-As a result, it remains an open question whether we can incorporate both properties.
-To this end, we propose a novel a novel Cross-experts Covariance Loss for Disentangled Learning model (Crocodile) framework, which we build this Covariance Loss (CovLoss) to alleviate embedding collapse while in the meantime enabling learning diverse user interests across various domains.
-To further unleash the power of Crocodile, we develop a novel Prior Informed element-wise Gating (PeG) mechanism to work with CovLoss.
-Empirical analysis on two public datasets confirms that Crocodile achieves state-of-the-art results.
-Further analysis also demonstrates that Crocodile can simultaneously mitigate embedding collapse and learn diverse representations.
-Ablation experiments reveal that each proposed component contributes positively to the overall performance.
-Our contributions can be briefly summarized as follows:
-  • We discover the mechanism of previous works in maintaining diversity, especially the trend of focusing more on bottom-level modules.
-We propose a novel metric, Diversity Index (DI), to quantify the diversity and conclude that the trend of previous works is optimizing to increase DI.
-  • We reveal that existing works fall into the dilemma of maintaining diversity and avoiding collapse, hence resulting in inferior performance.
-  • We propose the Crocodile approach, which successfully balances both properties.
-Empirical evidence on two public datasets demonstrates that Crocodile achieves state-of-the-art performance and resolves the aforementioned dilemma through comprehensive analysis.
-2. Related Work
-2.1. Multi-domain Recommendation
-Multi-domain learning (MDL) is a promising topic that has garnered attention from industrial recommendations.
-Many MDL methods have made great efforts to model complex multi-domain interests.
-Among them, MoE-based methods are widely applied in industrial recommendation systems, such as Taobao (Sheng et al., 2021) , Meituan (Chang et al., 2023) and Kuaishou (Zhou et al., 2023).
-STAR (Sheng et al., 2021) utilizes the dot product of domain-specific and domain-shared weight matrices to capture domain-specific interests.
-HiNet (Zhou et al., 2023) proposes two expert-like components to capture cross-domain correlations and within-domain information explicitly.
-PEPNet (Chang et al., 2023) introduces PPNet and EPNet, which incorporate domain information through gating.
-These methods often innovate from the perspective of making more bottom-level modules be domain-personalized within the model.
-Recently, STEM (Su et al., 2023) proposed a multi-embedding (ME) structure in multi-task learning (MTL) and step forward in allocating more bottom-level domain-personalized structure, i.e., embedding.
-It achieved better results compared to both single and multi-embedding versions of MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) in MTL experiments.
-However, STEM fails in the multi-domain recommendation.
-We will discuss the reasons in detail in the subsequent sections.
-Besides, some other paradigms are applied for multi-domain recommendation.
-AdaSparse (Yang et al., 2022) and ADL (Li et al., 2023) applied adaptive learning mechanism.
-EDDA (Ning et al., 2023a), COAST (Zhao et al., 2023) and UniCDR(Cao et al., 2023) provided graph learning-based solutions.
-2.2. Disentangled Representation Learning
-In the realm of disentangled representation learning (DRL), methods based on spatial or statistical relationships have been proposed.
-On the one hand, methods based on spatial relationships achieve disentangling by manipulating the distances or angles between representations.
-OLE (Lezama et al., 2018), dot product(Wu et al., 2021) and cosine similarity(Liu et al., 2021) push representations towards orthogonal.
-GAN-based models (Chen et al., 2016; Yu and Welch, 2021; Khrulkov et al., 2021) promote the discrimination among representations.
-However, they are computationally expensive and not conducive to online recommendation.
-On the other hand, statistical methods view representations as instances drawn from high-dimensional distributions and design loss based on statistical measures.
-Distance correlation (dCorr)(Liu et al., 2022a) encourages lower dCorr between different factors.
-ETL (Chen et al., 2023) models the joint distribution of user behaviors across domains with equivalent transformation.
-VAE-based methods (Ma et al., 2019; Wang et al., 2022; Liu et al., 2022b) encourage higher KL divergence between dimensions.
-However, they often assume a normal prior distribution, which is challenging to satisfy in the context of sparse interactions (Rendle, 2010), especially under sample volume imbalance of multi-domain recommendation.
-Besides, some methods integrate both aspects mentioned above.
-VICReg (Bardes et al., 2021; Shwartz-Ziv et al., 2024) employs covariance and variance as loss components within representations to promote independence across dimensions and prevent collapse while utilizing Euclidean distance between representations to ensure semantic similarity among representations of different patches of an image.
-2.3. Dimensional Collapse
-Dimensional collapse(Chi et al., 2022) is defined as representations that end up spanning a lower-dimensional space, which is evident through the pronounced imbalance in the distribution of singular values.
-Chi et al.
-(2022) found that SMoE structure (Fedus et al., 2022; Lepikhin et al., 2020) may lead to dimensional collapse due to the influence of gating.
-The work of (Guo et al., 2023b) claimed that this problem is also observed in popular models (e.g., DCN V2 (Wang et al., 2021) and Final MLP (Mao et al., 2023)).
-To measure the collapse, they proposed the Information Abundance (IA) (Guo et al., 2023b) metric, defined as the ratio between the sum and the maximum l1-norm value of the target singular values.
-A small IA indicates severe insignificance of tail dimensions and potential dimensional collapse.
-Beyond a single perspective, some methods (Hua et al., 2021) built the bridge between representation disentanglement and dimensional collapse prevention.
-It also claimed feature decorrelation helps alleviate dimensional collapse by reducing strong correlations between axes and promoting more diverse and informative representations.
+Further analysis reveals that Crocodile effectively mitigates embedding collapse while simultaneously learning diverse representations.
+Ablation experiments confirm the positive contributions of each proposed component to the overall performance.
+Online deployment on Tencent’s advertising platform shows a significant CTR and GMV lift.
+In summary, our contributions can be summarized as follows:
+  • We discover a dilemma between learning domain-specific characteristics and sufficient learning of parameters of existing multi-domain learning methods.
+  • We propose the Crocodile approach, which employs a Disentangled Multi-Embedding paradigm and a Prior Informed element-wise Gating (PeG) mechanism.
+  • Empirical evidence on two public datasets as well as online A/B testing on Tencent’s advertising platform demonstrate that Crocodile achieves state-of-the-art performance and resolves the aforementioned dilemma.
 2. On the Challenges of Multi-Domain Recommendation
-As we have discussed in Sec.
-1, existing works fall into the dilemma of avoiding dimensional collapse and maintaining interest diversity.
-In this section, we will discuss these two properties thoroughly.
-3.1. Preventing Dimensional Collapse
-According to Fig 1(a), domain-specific embeddings demonstrate its effectiveness in capturing diverse user interests.
-However, the naive adoption of domain-specific embeddings does not yield the expected outcomes.
-In fact, the methods with domain-specific embeddings, i.e., ME-PLE (Tang et al., 2020; Su et al., 2023) and STEM (Su et al., 2023), both performance deteriorate in the domain with less data.
-For example, in domain S6 of Kuairand1k (Gao et al., 2022) dataset, the ME-PLE and STEM show -0.4% and -0.1% lower AUC than the single domain method.
-This phenomenon has sparked our deep investigation into its underlying mechanism.
-Intuitively, the domain-specific embeddings are trained solely on data from the corresponding domain, which results in insufficient samples for domains with less data.
-Insufficient training data fails to provide adequate variance for high-dimensional representations, resulting in high-dimensional representations residing within a lower-dimensional subspace, called dimensional collapse (Guo et al., 2023b; Chi et al., 2022; Hua et al., 2021).
-To measure the collapse, we adopt the IA metric (Guo et al., 2023b), which is defined as $IA(E)=||\sigma||_{1}/||\sigma||_{\infty}$, where $E$ is the embedding of feature (e.g., item id) and $||\cdot||_{1},||\cdot||_{\infty}$ represent the $l1$-norm and $\infty$-norm, respectively.
-Remark 1.
-Intuitively, singular values $\sigma_{i}$ represent the significance of $i$-th singular value, and IA captures the balance of the singular value distribution.
-Thus, a small IA indicates severe insignificance of tail dimensions and potential dimension collapse.
+In this section, we study to what extent existing multi-domain recommendation models resolve the following two challenges: preserving each domain’s distinctiveness and sufficient learning of embeddings.
+Besides existing works, inspired by the success of multi-embedding (Su et al., 2023; Guo et al., 2023a), we developed two new baselines: a Shared and Domain-specific EMbeddings (SDEM) approach, which corresponds to the STEM in MTL, a Multi-Embedding PLE for multi-domain learning, ME-PLE in short, which corresponds to the ME-PLE in multi-task learning, and a Multi-Embedding MMoE for multi-domain learning, ME-MMoE in short.
+In the following, we evaluate the ability of existing and proposed methods to preserve each domain’s distinctiveness and the learning sufficiency of parameters.
+2.1. Preserving Distinctiveness
+Intuitively, the ability to capture various user interests across domains can be directly characterized by the diversity of representations.
+When facing conflict preference across domains, a set of diverse representations is enabled to encode both prefer/non-prefer semantics and thus provide distinct vectors to towers of those conflict domains.
+To this end, we propose a novel Diversity Index (DI) metric to quantify the diversity among the outputs of experts.
+We select a subset of samples where there is a diverse interest across domains, quantified by the output norm of domain-specific experts.
+We denote the set of samples satisfying such conditions as $\mathcal{D}_{(i,j)}$.
+Formally,
+(1) $\mathcal{D}_{(i,j)}=\{k|\|\bm{O}\|^{i}_{\text{SD}}(x_{k})\geq\tau_{t}\cap\|\bm{O}\|^{j}_{\text{SD}}(x_{k})\leq\tau_{b}\},$
+where $\tau_{t}$ and $\tau_{b}$ are two specified norm thresholds.
+Then, for those models where no experts are bound to any domains such as MMoE or STAR, we simply define the Diversity Index by whether there is one expert that learns a strong interest whilst another one learns a weak interest.
+That is, we just need to count the ratio of samples that have both a strong expert output norm and a weak one:
+(2) $\frac{\sum\limits_{k\in\mathcal{D}_{(i,j)}}\mathbb{I}[\exists p,q,(\|\mathbf{O}^{p}(x_{k})\|_{2}\geq\tau_{t})\cap(\|\mathbf{O}^{q}(x_{k})\|_{2}\leq\tau_{b})]}{|\mathcal{D}_{(i,j)}|},$
+where $\mathbb{I}[\text{{condition}}]$ is the indication function, which is assigned to 1 if the condition is true; otherwise, 0.
+$p,q$ denote the index of experts, $\mathbf{O}^{p}(x_{k})$ means the output of $p$-th expert of sample $x_{k}$.
+For those models with domain-specific experts, such as PLE, SDEM, and ME-PLE, we can quantify the Diversity Index by counting the ratio of samples that have a large expert output norm on domain $i$’s specific expert while a small norm on $j$’s expert:
+(3) $\frac{\sum\limits_{k\in\mathcal{D}_{(i,j)}}\mathbb{I}[(\|\mathbf{O}^{i}(x_{k})\|_{2}\geq\tau_{t})\cap(\|\mathbf{O}^{j}(x_{k})\|_{2}\leq\tau_{b})]}{|\mathcal{D}_{(i,j)}|}.$
+After we compute the diversity measurement $DI_{(i,j)}$ for all domain pairs $(i,j)$, we then take an average of them, getting the overall Diversity Index: $DI=\text{avg}(DI_{(i,j)})$.
+We then calculate the diversity index of several existing multi-domain recommendation models and those proposed models with multiple embeddings, as illustrated in Fig.
+1(a).
+We have the following findings.
+First, existing methods, in general, learn similar interests across domains, as indicated by the low diversity index.
+For example, the diversity index of HiNET, STAR, PEPNet, and PLE are less than 0.3.
+That means, for those samples that have shown strong domain-wise interest distinctiveness, these models can capture such distinctiveness for less than 30% of them.
+MMoE performs the best, with a diversity index of 0.615, but it’s still not a satisfactory result.
+Second, those multi-embedding models have higher diversity indices, especially compared to their single-embedding correspondence.
+ME-PLE gets a diversity index 0.403, compared to 0.284 of its single-embedding variant PLE; ME-MMoE gets a diversity index 0.716, compared to 0.615 of its single-embedding variant MMoE.
+SDEM gets the highest diversity index among models, i.e., 0.904.
+We conclude with the following finding:
+Finding 1.
+Existing single-embedding recommendation models for multi-domain learning fail to preserve domain distinctiveness.
+When employing multiple embeddings, especially domain-specific embeddings, models can preserve domain distinctiveness more decently.
+2.2. Sufficient Parameter Learning
+We found that the proposed ME-PLE and SDEM models perform poorly, particularly in domains with less data.
+For instance, in domain S6 of the Kuairand1k dataset (Gao et al., 2022), ME-PLE and SDEM show -0.4% and -0.1% lower AUC than the single domain method.
+We investigated the learned parameters, focusing on the embeddings.
+we assess the quality of embedding learning, particularly focusing on dimensional collapse testing, by analyzing the distribution of singular values and employing the aggregation metric Information Abundance (IA) (Guo et al., 2023b).
+It is defined as the sum of singular values divided by the largest one, i.e., $||\mathbf{\sigma}||_{1}/||\mathbf{\sigma}||_{\infty}$.
+A large IA indicates a balanced distribution of singular values and a robust embedding, while a small IA indicates an imbalanced distribution and dimensional collapse (Guo et al., 2023b; Chi et al., 2022).
 ((a)) $log(IA)$ of Experts
 ((b)) $log(IA)$ of Item Embeddings
-Figure 2.
-Information Abundance ($log(IA)$) dynamics of STEM’s bottom layer of experts and item embeddings on Kuairand1k.
-High and Low denote the items with the highest and lowest frequencies, respectively.
-Taking STEM as the example of methods applying domain-specific embedding, we analyzed its training IA dynamics of experts and embeddings in Fig.
+Figure 2. Information Abundance ($log(IA)$) dynamics of SDEM’s bottom layer of experts and item embeddings on Kuairand1k. High and Low denote the items with the highest and lowest frequencies.
+Taking SDEM as an example, we analyzed its training IA dynamics of experts and embeddings in Fig.
 2.
@@
 Within each domain, the low-exposure group consistently shows a lower IA than the high-exposure group.
-This further corroborates the notion that a lower exposure count (i.e., data volume) is more prone to dimensional collapse.
-Both analytical and empirical evidence suggest that inadequate data of domain can lead to dimensional collapse of the domain-specific embedding, hence leading to inferior performance.
-3.2. Interest Diversity
-Intuitively, the ability to capture various user interests across domains can be directly characterized by the diversity of representations.
-When facing conflict preference across domains, a set of diverse representations is enabled to encode both prefer/non-prefer semantics and thus provide distinct vectors to towers of those conflict domains.
-To this end, we propose a novel metric, the Diversity Index (DI), to quantify the representations’ ability to capture conflict preference across domains.
-Specifically, DI depicts the ratio of samples that contain both large and small norm of representations within the whole conflict samples, which is formulated as [^2] Here $AF$ and $SF$ are abbreviations for all forward and specific forward respectively, aligned with the concepts in  (Su et al., 2023).:
-$\mathcal{D}_{con(i,j)}$ contains preference conflict samples between domain $i$ and $j$.
-$\mathbb{I}[\text{{condition}}]$ is the indication function, which is assigned to 1 if the condition is true; otherwise, 0.
-$M$ represents experts, $\mathbf{O}_{k}^{(p)}$ means the output of $p$-th expert and $k$-th sample, $\tau_{t}$ and $\tau_{b}$ are the thresholds of big and small l2-norm.
-Remark 2.
-As shown in Fig.
-1(a), ME-MMoE (Guo et al., 2023a) exhibits significantly lower diversity compared to STEM (Su et al., 2023), and its improvement in diversity is relatively limited when compared to MMoE (Ma et al., 2018b).
-This suggests that the representations learned by ME-MMoE (Guo et al., 2023a) lack sufficient ability to solve preference conflict challenges.
+We conclude with the following finding:
+Finding 2. Models with domain-specific embeddings suffer from insufficient learning, indicated by dimensional collapse on embeddings of small domains, due to imbalanced data volume across domains.
 Figure 3. Crocodile Architecture, which consists of a Multi-Embedding (ME) layer, a Cross-expert Covariance Loss (CovLoss), and a Prior Informed Element-wise Gating (PEG) mechanism.
-4. Methodology
-We propose a multi-embedding architecture with Cross-experts Covariance Loss for Disentangled Learning (Crocodile) to resolve the dilemma, containing Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG), as shown in Fig.
+3. Our Method: Crocodile
+We propose a multi-embedding architecture with Cross-experts Covariance Loss for Disentangled Learning (Crocodile) to tackle the above-mentioned challenges.
+It consists of multiple embeddings and experts, with a disentanglement loss upon them, domain-specific towers, and a customized gating mechanism connecting the experts and the towers, as shown in Fig.
 3.
@@
 3.3. Cross-expert Covariance Loss
-Many pioneering work (Shwartz-Ziv et al., 2024; Bardes et al., 2021; Ermolov et al., 2021) have highlighted the efficiency of tailoring the covariance into objective functions.
-However, they aimed to decorrelate the dimensions within each representation rather than disentangle among representations of all experts.
-To learn diverse user interests at the expert level, we propose the cross-expert covariance loss (CovLoss) to explicitly disentangle representations among experts, defined as:
-(6) $\mathcal{L}_{Cov}=\frac{1}{d^{2}}\sum\limits_{p,q\in M\times M,p>q}||[\mathbf{O}^{(p)}-\overline{\mathbf{O}}^{(p)}]^{T}[\mathbf{O}^{(q)}-\overline{\mathbf{O}}^{(q)}]||_{1},$
+In order to make the output of experts uncorrelated so as to capture the user’s diverse or sometimes even conflicting interests on items, inspired by (Bardes et al., 2021; Yu et al., 2020), we propose a novel cross-expert covariance loss (CovLoss) to explicitly disentangle representations among experts, defined as:
+(7) $\mathcal{L}_{Cov}=\frac{1}{d^{2}}\sum\limits_{p,q\in M\times M,p\geq q}||[\mathbf{O}^{(p)}-\overline{\mathbf{O}}^{(p)}]^{T}[\mathbf{O}^{(q)}-\overline{\mathbf{O}}^{(q)}]||_{1},$
 where $\mathbf{O}^{(p)},\mathbf{O}^{(q)}\in\mathcal{R}^{N\times d}$ are the outputs of $p$-th and $q$-th experts, $\overline{\mathbf{O}}^{(p)}\in\mathcal{R}^{1\times d}$ is the average value of each dimension across all samples, and $||\cdot||_{1}$ is the $l1$-norm.
 Finally, the overall loss function is defined as :
-(7) $\mathcal{L}=\sum_{s=1}^{S}[\frac{1}{N_{s}}\sum_{k=1}^{N_{s}}\mathcal{L}_{BCE}(\hat{y}_{k}^{(s)},y_{k}^{(s)})]+\alpha\mathcal{L}_{Cov}$
+(8) $\mathcal{L}=\sum_{s=1}^{S}[\frac{1}{N_{s}}\sum_{k=1}^{N_{s}}\mathcal{L}_{BCE}(\hat{y}_{k}^{(s)},y_{k}^{(s)})]+\alpha\mathcal{L}_{Cov},$
 where $N_{s}$ is the number of $s$-th domain’s samples, $\hat{y}_{k}^{(s)}$ is the label of $k$-th sample in domain $s$, and $\alpha$ is the weight of CovLoss.
@@
 CovLoss aims to disentangle different representations by minimizing the covariance between dimensions among different experts.
-However, the existing gating structure hinders the effectiveness of CovLoss.
-MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020)’s gating mechanism assigns an identical weight to all dimensions within each expert; the dimensions within each expert are mutually constrained.
-PEPNet (Chang et al., 2023)’s proposed Gate Neaural Unit operates at the dimension level as well.
-However, it is designed to scale each dimension separately within the representation but not to disentangle these representations.
-Consequently, we propose an Element-wise Gating mechanism (EG), which implements weight control over corresponding dimensions between experts, as shown in Fig.
-3:
+However, the existing gating structure limits CovLoss’s effectiveness.
+The gating mechanisms in MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) assign identical weights to all dimensions within each expert, causing mutual constraints among dimensions.
+PEPNet’s (Chang et al., 2023) Gate Neural Unit operates at the dimension level, but it is designed to scale each dimension separately within the representation rather than to disentangle these representations.
+Consequently, we propose an Element-wise Gating mechanism (EG), which implements weight control over corresponding dimensions between experts, as shown in Fig. 3:
 (9) $t_{k}^{s}$ $=g^{s}_{k}\odot\mathbf{O}_{k},$
 where $\mathbf{O}_{k}\in\mathcal{R}^{K\times d}$ is the concatnation across all experts respected to $k$-th sample, and $\odot$ is the element-wise multiplication.
-Moreover, we are also concerned about the parameter burden of ME structure.
-In the original ME-MMoE structure, the number of embeddings is tied to the number of domains.
-This is because the gating is performed using $g^{s}_{k}=Softmax(e_{k}^{s}W_{g}^{s})$, which means the embedding $e_{k}^{s}$ is still related to a $s$-th gating.
-Thus, when it is deployed in situations with a large amount of domains, the scale of parameters becomes unmanageable.
+However, when the original ME is deployed in situations with a large number of domains, the scale of parameters becomes unmanageable, because the sets of embeddings are still related to the number of domains.
 To address this issue, we proposed Prior Informed Gating (PG), which involves an independent set of domain-unrelated prior embeddings $r_{k}$ to control the gating.
 $r_{k}$ takes the prior features as input, denoted as $x^{prior}$.
-For a given $k$-th sample and $f$-th field, the looked-up prior embedding is $e_{k,f}^{prior}$.
-Thus, the input of each gating is:
+For a given $k$-th sample and $f$-th field, the prior embedding is $e_{k,f}^{prior}$ and the input of each gating is:
 (10) $r_{k}=[e_{k,0}^{prior},e_{k,1}^{prior},\cdots,e_{k,F^{\prime}}^{prior}],$
@@
 where $r_{k}\in\mathcal{R}^{1\times l}$ is the prior embedding of $k$-th sample, $W_{g}^{s}\in\mathcal{R}^{l\times K\times d}$ and $Softmax(\cdot)$ is operated across the second dimension.
-Furthermore, PG can be understood as an attention (Vaswani et al., 2017)-like mechanism, comprising query, key, and value components corresponding to Prior embedding $r_{k}$, gating weight matrix $W_{g}^{s}$ of $s$-th domain, and expert outputs $\mathbf{O}_{k}$, respectively.
-Its significance lies in weighting the learnt user interests through the interaction of user, item, and domain information provided by query and key.
-4.5. Understanding the CovLoss
-CovLoss decorates the representations of different experts in an element-wise fashion.
-Beyond statistic understanding, CovLoss also directly addresses the core objective of cross-domain modeling: enhance the diversity of representations.
-For simplification, we consider the model involving only two experts, outputting $\mathbf{x}$ and $\mathbf{y}$ drawn from two high-dimensional distributions $\mathcal{X}$ and $\mathcal{Y}$, respectively, and the $i$-th element of $\mathbf{x}$ and $\mathbf{y}$ are $\mathbf{x}_{i}$ and $\mathbf{y}_{i}$, respectively.
-Following (Shwartz-Ziv et al., 2024), we assume the randomness comes from the input of the network.
-To maximize the diversity of the outputs from both experts, the objective for each pair of samples can be expressed as the expectation of l2-norm difference:
-(11) $\mathrm{E}[(||\mathbf{x}||_{2}$ $-||\mathbf{y}||_{2})^{2}]=\mathrm{E}[||\mathbf{x}||_{2}^{2}+||\mathbf{y}||_{2}^{2}-2||\mathbf{x}||_{2}||\mathbf{y}||_{2}]$
-$=\sum_{i=1}^{d}[\mu_{xi}^{2}+\sigma_{xi}^{2}]+\sum_{j=1}^{d}[\mu_{yj}^{2}+\sigma_{yj}^{2}]-2\mathrm{E}[\sqrt{\sum_{j\geq i}(x_{i}y_{j})^{2}}]$
-$\geq-2\mathrm{E}[\sqrt{\sum_{i=1}^{d}\sum_{j=1}^{d}(x_{i}y_{j})^{2}}]$
-$\geq-2\mathrm{E}[\sum_{i=1}^{d}\sum_{j=1}^{d}\sqrt{(x_{i}y_{j})^{2}}]=-2\sum_{j\geq i}\mathrm{E}[x_{i}y_{j}]$
-$=-2\sum_{i=1}^{d}\sum_{j=1}^{d}[\mu_{xi}\mu_{y_{j}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}cov(x_{i},y_{j})}],$
-where $x_{i},y_{j}\geq 0$ because the outputs from experts have been activated by ReLU before.
-According to the above Eq., minimizing $cov(x_{i},y_{j})$ can maximize the lower bound of diversity objective [^3] Although $\mu_{xi}\mu_{y_{j}}+cov(x_{i},y_{j})$ simultaneously incorporates both the first and second moments, they are nearly independent of each other in optimization.
+3.5. Discussion
+The Covariance Loss
+Many pioneering work (Shwartz-Ziv et al., 2024; Bardes et al., 2021; Ermolov et al., 2021) have highlighted the efficiency of tailoring the covariance into objective functions.
+Furthermore, some (Hua et al., 2021) claimed that feature de-correlation by standardizing the covariance matrix can help representations overcome dimensional collapse.
+These efforts have provided evidence to demonstrate the interpretability and efficacy of covariance.
+However, they aimed to de-correlate the dimensions within each representation rather than disentangle among representations of all experts.
+Besides, this loss shares the same spirit with the Maximal Coding Rate Reduction Principle (Chan et al., 2022; Yu et al., 2020) to de-correlate representations from different clusters (experts in our model).
+Figure 4. Architecture comparison of different MDL models with multiple embeddings.
+Connection to Existing Mulit-Embedding MTL or MDL Methods
+The proposed Crocodile differs from those existing multi-embedding-based methods in the following aspects: 1.
+Different from ME-PLE (Tang et al., 2020; Su et al., 2023) and SDEM (Su et al., 2023), which have domain-specific embeddings, all the embeddings and experts in Crocodile and ME-MMoE are shared across domains.
+Such design alleviates the insufficient learning of those embeddings corresponding to the small domains as illustrated in Sec.
+2.2.
+2.
+The embeddings and experts in Crocodile are disentangled with the Covariance Loss, making it able to preserve the distinctiveness of each domain, as illustrated in Sec.
+2.1.
+Fig.
+4 shows an architecture comparison of these multi-embedding MDL methods.
 4. Experiment
 In this section, we provide comprehensive analyses to address the primary research questions (RQs) outlined below:
-  • RQ1: How does the Crocodile perform compared with state-of-the-art recommenders in public datasets?
-  • RQ2: Can Crocodile capture diverse user interests without introducing dimensional collapse issue?
+  • RQ1: Can Crocodile outperform state-of-the-art methods on public datasets? Can this be attributed to maintaining domain distinctiveness while enabling sufficient representation learning?
+  • RQ2: Can CovLoss effectively disentangle the expert outputs?
   • RQ3: How does each proposed component contribute to performance gain?
-  • RQ4: How sensitive is Crocodile to critical hyper-parameters and model capacity?
-We employed two metrics to measure the model performance: AUC and gAUC. AUC measures the overall ranking performance, and gAUC evaluates the quality of intra-user item ranking.
+  • RQ4: How sensitive is Crocodile to hyper-parameters and model capacity?
+  • RQ5: How does Crocodile perform on a large-scale online recommendation system?
+The two public datasets correspond to video and product recommendations, respectively.
+The online revenue of these scenarios is often related to the ranking performance, ensuring that higher-scoring items better match user interests.
+Hence, we employed two metrics to measure the model performance: AUC and gAUC.
+AUC measures the overall ranking performance, and gAUC evaluates the quality of intra-user item ranking.
 4.1. Datasets
@@
 For Kuairand1k, we chose five domains with top-5 data volumes.
+Kuairand1k has 30 user-side and 62 item-side features.
 For AliCCP, we selected all three domains.
+AliCCP has 12 user-side, 4 item-side, and 4 user-item interaction features.
 Tab.
@@
 We removed features with fewer than 10 exposures and replaced them with default values.
+Both of these datasets exhibit issues of extreme data imbalance.
+For instance, in the Kuairand1k dataset, there is a 12-fold difference between the data in S0 and S6.
+Similarly, in the AliCCP dataset, S2 has 86 times larger data than S3.
 Table 1. Statistics of Kuairand1k and AliCCP datasets.
-#Sample #User #Item #Impr in Domain
-Kuairand1k (Gao et al., 2022) 11,648,743 1,000 188,148 2.4M(S0) / 7.8M(S1) / 0.4M(S2) / 0.9M(S4) / 0.2M(S6)
-AliCCP (Ma et al., 2018a) 85,316,519 241,080 499,321 32M(S1) / 52M(S2) / 0.6M(S3)
+#Sample #User #Item Domain (#Impr)
+Kuairand1k (Gao et al., 2022) 11,648,743 1,000 188,148
+S0(2.4M)/S1(7.8M)/S2(0.4M)
+S4(0.9M)/ S6(0.2M)
+AliCCP (Ma et al., 2018a) 85,316,519 241,080 499,321 S1(32M) / S2(52M) / S3(0.6M)
 4.2. Baselines
 To establish a performance benchmark for comparison, we will select comparative methods from both MTL and MDL for multi-domain recommendation.
-We apply the SharedBottom (Caruana, 1997) method in single/multi-domain contexts and subsequently introduce popular MTL methods such as MMoE (Ma et al., 2018b), PLE (Tang et al., 2020) (both Single-Embedding and Multi-Embedding (Su et al., 2023) versions), and STEM (Su et al., 2023).
+We apply the SharedBottom (Caruana, 1997) method in single/multi-domain contexts and subsequently introduce popular MTL methods such as MMoE (Ma et al., 2018b), PLE (Tang et al., 2020) (both Single-Embedding and Multi-Embedding (Su et al., 2023) versions).
+Inspired by STEM (Su et al., 2023), we also proposed the SDEM method with Shared and Domain-specific EMbeddings.
 Besides, we choose representative MDL methods, including STAR (Sheng et al., 2021), PEPNet (Chang et al., 2023), HiNet (Zhou et al., 2023), and AdaSparse (Yang et al., 2022).
-5.3. Performance Evaluation (RQ1 and RQ2)
+4.3. Performance Evaluation (RQ1)
 Table 2.
@@
 ME-PLE 0.73645 0.74240 0.80785 0.72727 0.75138 0.78499 0.61084 0.59717 0.55982 0.56170 0.61744 0.66075
-STEM 0.73741 0.74267 0.80526 0.72776 0.75449 0.78527 0.60820 0.59701 0.56250 0.56365 0.62595 0.66096
+SDEM 0.73741 0.74267 0.80526 0.72776 0.75449 0.78527 0.60820 0.59701 0.56250 0.56365 0.62595 0.66096
 MDL PEPNET 0.73732 0.74275 0.81396 0.72767 0.76239 0.78574 0.60812 0.59589 0.55435 0.55931 0.62733 0.66018
@@
 ME-PLE 0.62352 0.58727 0.62063 0.62132 0.58974 0.55777 0.59130 0.59049
-STEM 0.61978 0.58187 0.61978 0.61907 0.58941 0.55101 0.59132 0.59036
+SDEM 0.61978 0.58187 0.61978 0.61907 0.58941 0.55101 0.59132 0.59036
 MDL PEPNET 0.62273 0.59239 0.61988 0.62034 0.58851 0.56080 0.59085 0.58975
@@
 AdaSparse 0.62192 0.59656 0.62173 0.62189 0.58819 0.57170 0.58972 0.58924
-Crocodile 0.62448 0.59951 0.62192 0.62233* 0.59039 0.57085 0.59204* 0.59123*
+Crocodile 0.62485* 0.60033 0.62227 0.62327* 0.59062* 0.57237 0.59223* 0.59170*
 We compared Crocodile with state-of-the-art recommenders in the Kuairand1k (Gao et al., 2022) and AliCCP (Ma et al., 2018a) datasets.
@@
 4.3.1. Overall Performance.
-Our proposed Crocodile abrains the best overall AUC and gAUC in two datasets.
+Our proposed Crocodile achieves the best overall AUC and gAUC in two datasets.
 In the Kuairand1k dataset, Crocodile outperforms the second-best method significantly by 0.09% and 0.19% in terms of overall AUC and gAUC, respectively.
-In the AliCCP dataset, Crocodile surpasses the second-best method significantly by AUC of 0.04% and gAUC of 0.07% than the second-best method.
+In the AliCCP dataset, Crocodile surpasses the second-best method significantly by AUC of 0.138% and gAUC of 0.12% than the second-best method.
+Usually, a 0.1% AUC (gAUC) lift is regarded as a huge improvement in recommendation systems (Zhu et al., 2022).
 Crocodile achieves larger AUC and gAUC almost for almost all domains in two datasets significantly compared to single domain method, with p-value¡0.05.
-Such positive transfer can be attributed to Crocodile’s better usage of multi-domain information and multi-embedding to enhance performance in each domain, without raising dimensional collapse.
+Such positive transfer can be attributed to Crocodile’s preserving domain distinctive and sufficiently representation learning.
 Figure 5.
 Comparison of the singular value spectrum $log(\sigma)$ of item ID and user ID embeddings in Kuairand1k dataset.
-We reported $log(\sigma)$ of ME-PLE and STEM S6-specific embedding, while the original $log(\sigma)$ of single embedding or the average of other ME methods.
-5.3.2. Comparison with MTL for MDR methods.
-Multi-task learning (MTL) methods for Multi-Domain Recommendation (MDR) can be primarily categorized into single-embedding (SE) and multi-embedding (ME) paradigms (Su et al., 2023).
-We observed that some methods allocating domain-specific embedding underperform compared to SE methods, despite the fact they can capture diverse user interests in Fig.
-5(a).
-For instance, in the comparison between PLE and ME-PLE on the Kuairand1k dataset in, Tab.
-2, PLE achieved an overall AUC of $0.78565$ and gAUC of $0.66093$.
-In contrast, ME-PLE’s AUC and gAUC were $0.78499$ and $0.66075$, respectively, both lower than those of PLE.
-To investigate the underlying reason, we conducted a domain-by-domain performance comparison and found that ME-PLE exhibited notably poorer results in the smallest domain, S6, with a negative transfer in AUC of -0.4% compared to the single domain.
-A similar negative transfer was also observed in STEM.
-We analyzed their embeddings via singular spectrum analysis (Guo et al., 2023b) and show the results in Fig.
-4.
-We discovered that ME-PLE and STEM, in comparison to other methods, exhibited a clear dominance of top factors with much higher singular value over tail factors in the S6 embeddings, indicating the dimensional collapse issue.
+We reported $log(\sigma)$ of ME-PLE and SDEM S6-specific embedding, while the original $log(\sigma)$ of single embedding or the average of other ME methods.
+4.3.2. MDL methods with domain-specific embeddings.
+The Mulit-Embedding PLE (ME-PLE) and SDEM are two new variants with domain-specific embeddings, inspired by task-specific embeddings in multi-task learning.
+However, ME-PLE’s overall AUC and gAUC are lower by 0.07% and 0.02% compared to PLE, respectively.
+Analyzing each domain, we found the smallest domain experienced the most significant drops in AUC and gAUC, both exceeding 0.3%.
+A similar negative transfer was also observed in SDEM.
+To further investigate the mechanism, we examined the quality of representation learning by observing the imbalance in the distribution of singular values (Guo et al., 2023b; Chi et al., 2022).
+As shown in Fig.
+5, We discovered that ME-PLE and SDEM exhibited a clear dominance of top factors with much higher singular values over tail factors in the S6 embeddings.
+This means that their embeddings of small domains suffer from insufficient learning.
 In contrast, Crocodile’s embeddings exhibit a more balanced importance of singular values, thereby better supporting the high-dimensional representational space.
-Consequently, within domain S6, Crocodile outperforms ME-PLE by 0.68% in AUC and 1.8% in gAUC, and surpasses STEM by 0.37% in AUC and 0.98% in gAUC.
-Even though STEM learns the most diverse user interest, it still fails to surpass Crocodile due to dimensional collapse.
+Consequently, within the smallest domain S6, Crocodile outperforms ME-PLE by 0.68% in AUC and 1.8% in gAUC, and surpasses SDEM by 0.37% in AUC and 0.98% in gAUC.
+Even though SDEM learns the most diverse user interests, it still fails to surpass Crocodile due to insufficient representation learning.
 ((a)) Crocodile vs MDL and MTL
@@
 We compared Crocodile with MDL and MTL methods in (a) and other disentanglement losses in (b).
-5.3.3. Comparison with MDL methods.
-The Multi-domain Learning (MDL) approaches typically evolve on a trend of progressively allocating domain-personalized components from top to bottom to capture more diverse user interest.
-To empirically support this, we analyze the Diversity Index (DI) of both MDL and MTL methods in Fig.
-5(a).
-Firstly, along with the trend from top to bottom, the DI of STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), and STEM (Su et al., 2023) reveals an increased trend.
-Secondly, methods with an ME structure typically exhibit higher DI than their SE versions, e.g., ME-PLE has higher DI than PLE.
-However, as discussed in Section 5.3.2, the allocation of domain-specific embeddings can lead to the side effect of dimensional collapse, which in turn may result in a negative transfer.
-Consequently, we are curious as to whether methods that do not employ domain-specific embeddings can achieve favorable outcomes.
-For example, HiNet (Zhou et al., 2023) only allocates domain-specific components at the expert level and displayed the best ability to prevent dimensional collapse in Fig.
-4.
-It thus got the best AUC of 0.76451 in domain S6, in Kuairand1k.
-However, HiNet still falls in suboptimal performance in overall AUC and gAUC in both datasets compared to Crocodile.
-This is because of its low ability to capture diverse user interests in  5(a).
-Conversely, Crocodile achieves the second-best DI, only slightly trailing behind STEM, which employs domain-specific embeddings but faces the issue of dimensional collapse.
-Consequently, Crocodile attains the best overall AUC and gAUC.
+((a)) CovLoss between Experts
+((b)) Covariance across all dimensions of all Experts
+Figure 7.
+Heatmaps of CovLoss between Experts (left) and Covariance across all dimensions of all Experts (right) on Kuairand1k.
+In the comparison methods with the same structure, ”without CovLoss” corresponds to the loss function being only BCE Loss, while ”with CovLoss” corresponds to the loss function as the combination of BCE Loss and CovLoss.
+4.3.3. Existing MDL methods.
+Multi-domain learning (MDL) methods also exhibit a top-down structural evolution, explicitly capturing domain-specific interests by allocating domain-specific components layer by layer.
+We expect that the more thorough the design, the more domain-distinctive the learnable parameters, resulting in better performance.
+We use the IA metric introduced in Sec.
+2 to quantify it, as shown in Fig.
+6(a).
+Along with the trend from top to bottom, the DI of STAR (Sheng et al., 2021), HiNet (Zhou et al., 2023), PEPNet (Chang et al., 2023), and SDEM (Su et al., 2023) reveals an increasing trend, which is consistent with expectations.
+However, the performance does not follow this order.
+For example, SDEM has a better DI than HiNet but exhibits lower AUC and gAUC, because performance is also influenced by the sufficiency of representation learning, as discussed in Sec.
+2.2.
+In contrast, Crocodile maintains domain distinctiveness more effectively while ensuring adequate representation learning.
+Although HiNet exhibits more balanced singular values compared to Crocodile, it performs poorly in maintaining domain distinction.
+Consequently, Crocodile achieves higher AUC and gAUC than HiNet by 0.12% and 0.24%, respectively.
 4.3.4. Comparison with Other Forms of Losses.
-Through comparison with MTL and MDL approaches, we found that Crocodile captures the diverse interests of users (Fig.
-5(b)) without introducing the issue of dimensional collapse (Fig.
-4), thereby achieving the best overall AUC and gAUC scores in Kuairand1k and AliCCP.
-Thus, can other disentangled losses also achieve favorable diversity and corresponding positive outcomes compared to those without them?
-We compare CovLoss with several other disentangling losses, including Dot Product (Dot) (Wu et al., 2021), Cosine Similarity (Cos) (Liu et al., 2021), Distance Correlation (dCorr) (Liu et al., 2022a) and OLE (Lezama et al., 2018) (F-norm form).
-We first analyzed the ability to capture user diverse interests in Fig.
-5(b).
-We observed that all disentangled losses could achieve higher DI than ME-MMoE-PeG, where ME-MMoE-PEG is the basic structure without any disentangled loss compared to others.
-Table 4.
-Comparison of different disentanglement losses with the base model on overall AUC and gAUC in Kuairand1k.
-* indicates that the performance difference against base is significant at p-value¡0.05.
+In Sec.
+4.3.2 and  4.3.3, we found that Crocodile can achieve sufficient representation learning while maintaining domain distinctiveness, resulting in the best performance on two public datasets.
+This raises the question: can other loss functions achieve similar benefits?
+We conducted experiments by replacing the loss function in Crocodile, naming the method that retains the structure but uses only BCE Loss as Base.
+Table 4. Comparison of different losses with the base model on overall AUC and gAUC in Kuairand1k. * indicates that the performance difference against base is significant at p-value¡0.05.
 Losses AUC Lift(%) gAUC Lift(%)
 Base 0.78611 - 0.66208 -
-+Dot (Wu et al., 2021) 0.78667 0.06% 0.66285 0.08%
++Dot (Wu et al., 2021; Tian et al., 2023) 0.78667 0.06% 0.66285 0.08%
 +Cos (Liu et al., 2021) 0.78634 0.02% 0.66263 0.06%
@@
 +CovLoss
-0.78683 0.07% 0.66373 0.17%
-Among them, we found that dCorr (Liu et al., 2022a) loss achieves the smallest improvement of only 0.01% and 0.01% for AUC and gAUC lift, respectively in Tab.
-4.
-This is coherent with its related poor diversity compared to other disentangled losses in Fig.
-5(b).
-In contrast, ultilizing CovLoss achieves the best AUC and gAUC with the lift of 0.07% and 0.17% compared to the Base model.
-This evidence supports that the ability to learn diverse interests is important to improve performance.
-Furthermore, we are interested in exploring whether the transfer of other types of loss functions, e.g., load balance (Shazeer et al., 2017) and transfer learning (Chen et al., 2023), could yield performance enhancements.
+0.78683* 0.07% 0.66373* 0.17%
+Among the loss functions in Tab.
+4, dCorr (Liu et al., 2022a) shows the smallest improvement, with only a 0.01% increase in AUC and gAUC, consistent with its poor domain distinctiveness in Fig.
+6(b).
+In contrast, CovLoss achieves the highest AUC and gAUC improvements of 0.07% and 0.17% over the Base model, highlighting the importance of domain distinctiveness for performance enhancement.
+Additionally, we explored other loss functions, such as load balance (Shazeer et al., 2017) and transfer learning (Chen et al., 2023), for potential performance gains.
 As shown in Tab.
-4, the Importance loss (Shazeer et al., 2017) resulted in a limited improvement, whereas Trans5, one of the effective versions of equivalent transformation losses in (Chen et al., 2023), had a detrimental effect.
-This can be attributed to that these loss functions do not explicitly promote diversity in learning.
-Instead, they focus on balancing the contributions of various experts or enhancing the robustness of representations, which does not meet the expected outcomes.
+4, Importance loss (Shazeer et al., 2017) offered limited improvement, while Trans5, an effective transformation loss from (Chen et al., 2023), had a negative impact.
+They all fail to surpass CovLoss.
+4.4. CovLoss Analysis (RQ2)
+We are curious about if the CovLoss truly disentangled the experts as expected.
+We conducted experiments using the same backbone on Kuairand1k, with and without CovLoss in Fig.
+7.
+We progressively reveal the effectiveness of CovLoss from the course, i.e., expert level, to fine, i.e., dimension level.
+Firstly, at the expert level in Fig.
+7(a), the method with CovLoss shows significantly lower CovLoss values across each pair of experts compared to the baseline.
+For example, the CovLoss between the 3-th and 5-th Experts in the method with CovLoss is $3\times 10^{3}$ lower than that without CovLoss.
+Secondly, at the dimension level in Fig.
+7(b), the dimensional-wise heatmap of the method with CovLoss has much smaller number of high-value points than without CovLoss, shows it also has overall significantly smaller covariance compared to the baseline.
+Please note that CovLoss acts as an auxiliary loss, serving as a soft constraint and certain learnable parameters must retain intensity information.
+Thus, few high-value points are still observed in the method with CovLoss.
+Finally, in terms of experimental performance, the method with CovLoss achieved AUC and gAUC significant improvements of 0.07% and 0.17%, respectively, on the Kuairand1k dataset compared to the method without CovLoss, as shown in Tab.
+5.
+These findings demonstrate that CovLoss effectively disentangles the experts and enhances model performance.
 4.5. Ablation Study (RQ3)
@@
 ME PG B+C 0.78642 0.05% 0.66333 0.28% 0.62214 0.28% 0.59124 0.13%
-Crocodile ME PEG B+C 0.78683* 0.09% 0.66373* 0.32% 0.62237* 0.30% 0.59134* 0.14%
+Crocodile ME PEG B+C 0.78683* 0.09% 0.66373* 0.32% 0.62327* 0.49% 0.59170* 0.18%
 In this section, we empirically test the effectiveness of our proposed components with an ablation study.
@@
 In particular, CovLoss is more critical in the Kuairand1k dataset, while CovLoss and EG are equally important in AliCCP.
-Table 6. gAUC metric along with different numbers of embeddings (#Emb) on ME-MMoE-PG and Crocodile. * indicates that the performance difference against ME-MMoE-PG is significant at p-value¡0.05.
-#Emb 2 3 4 5 6 7 8 9 10
-ME-MMoE-PG 0.66138 0.66157 0.66206 0.66232 0.66203 0.66196 0.66286 0.66279 0.66285
-Crocodile 0.66198* 0.66284* 0.66324* 0.66373* 0.6637* 0.66394* 0.66395* 0.66432* 0.66397*
-Lift (%) 0.06% 0.13% 0.12% 0.14% 0.17% 0.20% 0.11% 0.15% 0.11%
-5.5. Hyper-Parameters and Embedding Sets (RQ4)
-In this section, we investigate the sensitivity of the model’s performance to hyperparameters and the number of embeddings.
-For hyperparameters, the primary parameter is the weight of CovLoss in Eq.
-7, i.e., $\alpha$.
-Taking Kuairand1k as an example, we find that as long as $\alpha$ is not too small (below $2\times 10^{-5}$), Crocodile’s AUC and gAUC remain ¿0.78590 and ¿0.66174, thus maintaining its status as the second-best method.
+4.6. Setting Sensitivity Analysis (RQ4)
+In this section, we investigate the sensitivity of the model’s performance to hyper-parameters, for example, the weight of CovLoss $\alpha$ and the number of embeddings.
+At Kuairand1k, we find that as long as $\alpha$ is not too small (below $2\times 10^{-5}$), Crocodile’s AUC and gAUC remain ¿0.78590 and ¿0.66174, thus maintaining its status as the second-best method.
+Figure 8. gAUC metric of Crocodile varying different numbers of embeddings (#Emb) compared with Base, i.e., ME-MMoE with PG, in Kuairand1k. The solid line represents the lift ratio.
 Additionally, we analyzed the impact of model capacity on performance by adjusting the number of embedding sets.
-At the minimum number of MEs (i.e., 2-sets), Crocodile achieves a gAUC of 0.66198 and an AUC of 0.78595, which still surpasses other MTL and MDL methods, even when the MTL(ME) methods apply for 5-sets of embeddings.
-Moreover, as the number of embeddings increases, the performance gains of Crocodile over ME-MMoE-PG are robust ($+0.06\%\sim+0.20\%$), as illustrated in Tab.
-6.
-5.6. Computational Complexity
+At the minimum number of MEs (i.e., 2-sets), Crocodile achieves a gAUC of 0.66198 and an AUC of 0.78595, which still surpasses other MTL and MDL methods, even when the other MTL(ME) methods employ 5 embeddings sets.
+Moreover, as the number of embeddings increases, the performance gains of Crocodile over ME-MMoE with PG are robust ($+0.06\%\sim+0.20\%$).
+The detailed results are shown in Fig.
+8.
+4.7. Online Deployment (RQ5)
+We conduct an empirical evaluation of Crocodile on Tencent’s online advertising platform.
+The underlying model architecture utilizes Heterogeneous Experts with Multi-Embedding framework (Pan et al., 2024).
+Specifically, our approach involves training multiple feature interaction experts, such as GwPFM (Pan et al., 2024) (a variant of FFM (Juan et al., 2016) and FwFM (Pan et al., 2018)), IPNN (Qu et al., 2016), DCN V2 (Wang et al., 2021), or FlatDNN, to capture diverse feature interactions for sparse ID features.
+Additionally, we learn multiple embedding tables for all features (Guo et al., 2023b; Su et al., 2023; Pan et al., 2018), with each table corresponding to one or several experts.
+We handle sequence features by TIN (Zhou et al., 2024) to capture the semantic-temporal correlation, numeric features by the multiple Numeral Systems Encoding (Chen et al., 2022; Pan et al., 2024), and embedding features from external models via Similarity Encoding Embedding (Pan et al., 2024).
+Specifically, we train a multi-domain model over two main scenarios, with two embedding tables and experts. The output of these two experts is then fed into the primary scenario.
+We conduct online A/B testing on March 2024.
+The baseline model employs a common Embedding & Interaction architecture.
+During 5% A/B test, we observe a 0.11% AUC lift on several main tasks such as convertible click prediction, and a significant 0.72% CTR lift and 0.73% GMV lift.
+False Tolerence
+During online training, if the training samples from the auxiliary domains become unavailable or have some issues, we simply roll back the model to the latest valid checkpoint and keep it trained and updated based on only the main domain until the auxiliary one is recovered.
+If the main domain’s data pipeline has any problems, we roll back the model and keep it fixed.
+5. Related Work
+5.1. Multi-domain Recommendation
+Multi-domain learning (MDL) is a promising topic that has garnered attention from large-scale industrial recommendations, such as Alibaba (Sheng et al., 2021; Zhang et al., 2022; Tian et al., 2023), Meituan (Chang et al., 2023) and Kuaishou (Zhou et al., 2023).
+Many MDL methods (Sheng et al., 2021; Zhou et al., 2023; Zhang et al., 2022; Li et al., 2023a; Chang et al., 2023; Tian et al., 2023; Wang et al., 2024) have made great efforts to model complex multi-domain interests.
+Current methods often start from the Shared-bottom (Caruana, 1997) structure to explore capturing both domain-shared and domain-specific interests.
+STAR (Sheng et al., 2021) introduces a learnable matrix and dot product operator before outputting the prediction value.
+HiNet (Zhou et al., 2023) and HAMUR (Li et al., 2023a) allocate domain-shared and domain-specific networks.
+Furthermore, the PEPNet (Chang et al., 2023) method introduces EPNet, which performs dot product operations between domain-related and domain-shared embeddings, while Maria (Tian et al., 2023) designs feature scaling, refinement, and correlation modeling to capture cross-domain interests at the feature-level.
+They show the trend of integrating domain information at the more bottom level, i.e., from Tower to Embedding.
+Recently, STEM (Su et al., 2023) proposed a multi-embedding (ME) structure in multi-task learning (MTL) and step forward in allocating domain-personalized structures.
+It achieved better results compared to both single and multi-embedding versions of MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) in MTL experiments.
+However, STEM fails in the multi-domain recommendation.
+We will discuss the reasons in Sec.
+2.
+Besides, some other paradigms are applied for multi-domain recommendation.
+For example, AdaSparse (Yang et al., 2022) and ADL (Li et al., 2023b) applied adaptive learning mechanism.
+M2M (Zhang et al., 2022) utilizes meta-learning techniques.
+EDDA (Ning et al., 2023a), COAST (Zhao et al., 2023) and UniCDR(Cao et al., 2023) provided graph learning-based solutions.
+5.2. Disentangled Representation Learning
+Disentangled representation learning helps preserve the distinctiveness of embeddings.
+Loss-based methods such as OLE (Lezama et al., 2018), dot product (Wu et al., 2021), and cosine similarity (Liu et al., 2021) push representations of different domains towards orthogonality and residing in corresponding subspace.
+Distance correlation (dCorr) (Liu et al., 2022a) encourages lower dCorr between different factors.
+ETL (Chen et al., 2023) models the joint distribution of user behaviors across domains with equivalent transformation.
+VAE-based methods (Ma et al., 2019; Wang et al., 2022; Liu et al., 2022b) promote higher KL divergence between dimensions but often assume a normal prior distribution, which is challenging in the context of sparse interactions (Rendle, 2010), especially with imbalanced sample volumes in multi-domain recommendation.
+GAN-based models (Chen et al., 2016; Yu and Welch, 2021; Khrulkov et al., 2021) enhance discrimination among representations but are computationally expensive and unsuitable for online recommendation.
+Recently, some research has provided systematic and unified insights.
+VICReg (Bardes et al., 2021; Shwartz-Ziv et al., 2024) uses covariance and variance as loss components to promote independence across dimensions while preventing insufficient learning.
+Methods like (Chan et al., 2022; Yu et al., 2020) propose a maximal coding rate reduction principle on mixed distributions to maximize the coding rate of the entire representation while minimizing it within each cluster.
+5.3. Dimensional Collapse
+Dimensional collapse (Chi et al., 2022; Guo et al., 2023b) occurs when all representations span a lower-dimensional space, reflecting the insufficient learning of embeddings.
+Chi et al.
+(2022) found that the SMoE structure (Fedus et al., 2022; Lepikhin et al., 2020) may cause dimensional collapse due to gating influences.
+(Guo et al., 2023b) observed this issue in popular models such as DCN V2 (Wang et al., 2021) and Final MLP (Mao et al., 2023).
+To measure collapse, they proposed the Information Abundance (IA) (Guo et al., 2023b; Pan et al., 2024) metric, defined as the ratio between the sum and the maximum l1-norm value of the target singular values.
+A small IA indicates severe insignificance of tail dimensions and potential dimensional collapse.
+Some methods (Hua et al., 2021) bridge the gap between representation disentanglement and dimensional collapse prevention.
+They also claim that feature decorrelation helps alleviate dimensional collapse by reducing strong correlations between axes, promoting more diverse and informative representations.
+6. Conclusion
+Multi-domain learning is vital in enhancing personalized recommendations.
+Existing methods fail to learn both dimensional-robust representations and diverse user interests.
+We propose Crocodile, which employs a Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG) based on the multi-embedding (ME) paradigm.
+Both offline and online experiments validate the effectiveness of Crocodile.
+Appendix A Computational Complexity
 The Multi-Embedding (ME) paradigm has already been widely adopted in industry (Zhaok et al., 2021; Ning et al., 2023b; Su et al., 2023; Guo et al., 2023b; Pan et al., 2024).
 Compared to these ME architectures, our proposed Crocodile introduces the CovLoss upon it, which only influences the training process while not the inference at all.
-Its complexity is not only small and negligible when compared with the ME-based backbone, but it can be further reduced by sampling.
+Its complexity is not only small and negligible based on the ME-based backbone, but it can be further reduced by only computing the loss based on downsampled samples.
 The complexity of CovLoss is $\frac{M(M-1)}{2}\times(d^{2}N)$, shown lower complexity than dCorr (Liu et al., 2022a), which is $\frac{M(M-1)}{2}\times(d^{3}N+d^{2}N)$, where $M$ and $N$ are the number of experts and samples, and $d$ is the dimension of expert output.
@@
 We discovered that selecting as low as 32 samples yields no significant difference in computed covariance norms (p-value¿ 0.1) in the testing dataset, contributing as much as 99.2% complexity reducing, as we originally selected $N=1024$.
-6. Conclusion
-Multi-domain learning is vital in enhancing personalized recommendations.
-Existing MTL methods fail to learn both dimensional-robust representations and diverse user interests.
-We propose Crocodile, which employs a Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG) based on the multi-embedding (ME) paradigm.
-Comprehensive experimental evaluations validate the effectiveness of Crocodile.